A Bibliography of Publications in Science and Public Policy: 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 0.01

Title word cross-reference

2 [Gri84].

12.95 [Col83]. 17.50 [Ano84a]. 1802-03 [Bro84]. 1845 [Hel83b]. 1950s [Uen86]. 1970s [Tal80]. 1980 [Jon81a]. 1980s [Ano80i, Cer80, Hel81c, Wie81]. 1981 [DB82]. 1982 [Kru84a]. 1983 [Por84a]. 1984 [Por84b]. 1990s [Ano80g].

21st [Hou87, Nix87]. 225pp. [Dav84].

4.95 [O’T83].

= [Kel82]. abolition [Bra81a]. absurdities [HR85]. Academic [Bro83a, Rog81, Cla88, Irw89, ZL87]. Academy [Baa88]. Acceptability [Gri80, GB80]. access [Man82]. accessible [Har88b]. accountability [Bra80]. achieving [Haq89b]. Acid [DB82, Ash82]. activities [Gia86c, Hoo86]. ad [Ano80a]. Adams [G.84]. Added [Bra81a]. administration [Fam86]. advice [Fre88, Tal89, O’T83]. Advisory [Ano83e]. Advocacy [BC82a]. affluence [For86d]. Africa [EDR85, For85c, For85b, For86c, Jon85b, pJAH83b, Man80a, Maz80, Nde85, Tit85, Wad85]. African [Tai83a]. Afro [Moo86]. Afro-centric [Moo86]. After [Ber89, GM89a, BC82a]. again [Har83]. against [Dav82, For88a]. age [Cam82, Cro84, Gia86b, Hil84c, Lal84, Str84]. agenda [Lei88]. agendas [Ein84]. Agricultural [Bal88, Fam86, Cox83, Rea89]. Agriculture [KM89a, Ega88, RDU87]. Ahead [Smi89, Fel81]. aid [Tai83c]. air [JS83]. aircraft [Zuc81]. Alan [Hil84a]. Aldershot [Dav84, Hil84a]. aliens [Bak87a]. alone [Kec84]. alter [BS83c]. Alternative [Rob81a, Cro89a, Ste82]. Alternatives [RCH87, Swe87]. Alvey [Oak84]. Always [dlM88a]. Amateur [Mea86]. Amazing [Wat83c]. America [Hel83b, Ash83, Wat82c, Wat83e, Wat83f]. American [BS83f, Col88b, DeW82, Tor84]. Amsterdam [Blu85, ZL87]. analyses [Ash82]. analysis [Ano88g, BM88a, LLB86, Sha88a, Sha88b, Wat83a]. ancient [Pyk87]. Anglophone [EDR85]. Animal [Sco89]. Annex [Rot81]. anniversary [Gol84e]. Announcement [Mg86]. Anomalies [HR85]. anthropology [Sap89]. Anti [Bra84, Ano84b]. Anti-gravity [Bra84, Ano84b]. apathy [dlM88d]. application [JS83]. Applications [Hes82a, Sim80]. Applied [Irw86, Rea89]. appraisal [Bro83g]. approach [Car86, Chu87, Per85, Ray86]. Approaches [(Re84, Lic87, Mor81c, NSA^{+}88, Yox87]. appropriate [Cro89a]. Arab [BC82a, Gor84, Sab80]. Arabian [BC82a]. architecture [GD84]. archives [RT84]. areas [NF82]. arguments [Wat83e]. arid [RDU87]. armament [BV81]. arms [Bel86, Blo83, Col84, DeW82, Gal87]. arousal [Don84]. Artificial [Arn85, Jah86]. Arts [Moo89]. ASEAN [Ano84e, Ghe82]. ASEAN-EEC [Ano84e]. Ashby [Ash82]. Asia [Hil86a]. Asia/Pacific [Hil86a]. aspect [Boy83]. Aspects [Hil84b, Jon85b, aH81]. aspirations [Ano83e]. Assessing [Gut80, RF80]. Assessment [RD80, Car86, Coo84, JS83, TS83c]. assets [Bud89a]. assistance [Ega88, Mor81a].

Assistants [Dav84]. Association [(Re84]. assumptions [Wie81]. AT&T [Por84b]. attacks [BC82a]. attainable [Bel86]. Attitudes [Has84, BFSL85, KW87, Wei89]. Attracting [Oke86]. Australia [Dod89b, Jos88, Jos89]. Australian [Low83, Sco89]. authoritative [Zim89]. authority [Smi86]. Authors [Ano84g, Ano84h, Ano87m, Ano87n]. autobiographies [Sho88]. Autobiography [Gol81b]. automated [Ros83]. automaticity [Don84]. Ayangba [Ega88].

B [Dav84, O’T83]. back [Gum85, Neu86]. Baedeker [BC83g]. baked [Ano85j]. Bakewell [Dav84]. balanced [Dod89a]. Bank [Ega88]. Bap [Gum87a]. Barons [Hes82a]. Barriers [Lan89]. base [Ram88b]. based [Few80, Oke86, RD89]. Basic [Jor86]. basin [Sah80, Tho82]. Basis [Sha88a]. Basque [Iri85]. Bay [Sut80]. be [Ros83]. belated [TS83c]. Belgrade [Emm80]. Bell [Por84b]. benefits [Low83, Mor85]. Bep [Gum87a]. Bernal [Ash80, GM89b, Rav82]. best [Irw86]. better [Duc83]. between [BC80b, For86d]. Beyond [Bak83, Wal83a]. Bibliography [Ano82g, Ano85a, Ano86a]. Bibliometric [Nar87]. Big [Ada82]. biomass [Few80]. biomass-based [Few80]. Biotechnology [Här88a, Yox87, Cum87]. bird [Tit85]. birth [Div88, Mep89]. birthright [Mep89]. black [Mor81b, Tit85]. bland [Boy88]. board [Ano80q]. Bomb [Bak83, Mey82]. Book [Ano84a, Col83, Dav84, Hil84a, Jon81a, Kru84a, O’T83, Por84a, Por84b, Syr83]. Booknotes [For85a]. Born [Mey82]. Botany [Bel82]. bound [Ano84k]. boundary [Wor89]. Braefoot [Sut80]. Brandt [Har83, Mor80, Tra83a]. Brazier [Ano84b]. Brazil [Wat83d]. Brazilian [BS84b, dMSG89, RCH87, Tig83]. Bread [Bar84]. breastfeeding [Mep89]. breeding [Web89]. Brett [BC83g]. Brett-Crowther [BC83g]. brewing [Mar87b]. bridge [Ano80c]. Britain [Bon86a, FGSA86, G.85a, Gre82a, Gre83, Jac87, Köö85, Thr83b]. British [BC82a, BS83h, Bro84, Gar87, Gre87, Jas89, Ley88a, Mat83, Mor86, Pal87]. Britten [Bri82]. Broading [Yox85a]. Brown [G.85b]. budworm [BC81e]. Building [BC82a, Tai83b]. business [Hou83, Van87].

C [Ano84a]. cabin [Zuc81]. calculus [HL88]. camp [Coh81]. Campaign [For88a, Por84a]. Can [Kel82, Hay89b]. Canada [GR88, HB88, Lei88, Wil80, Wil82]. Canadian [DG88, Ste84, Wil89a, Wil89b, dlM88b]. cancer [BBGM85, Dha84, (Re84, Rob81b, (Re84]. capital [SW86]. car [Hil85]. Carver [Bul83]. Case [Bis80, Kib81, Mey82, BM88b, Bro83g, Jac87, Lei88, Ram88b, Sta81, Web89]. case-study [Sta81]. cases [Hil86a]. Catastrophe [Bow84b]. cautionary [NSA^{+}88]. CEGB [Ash82]. centrally [Jas89]. Centre [Fro84, Joh86, Lak81]. centred [Smi87]. centric [Moo86]. century [Han84, Hou87, Nix87]. chain [Bar83a]. Challenge [Gia86a, Mor86, Mor81a, Nar88, Pon88]. Change [Gol82, Joh86, Dav87b, Hel81d, Ing87, Jac80, Jen82a, Man80b, Mor81b, pAM81, Mus80, Tal84, WS83]. changes [Bow84b]. Changing [Ale85, Mag84, Spe81, dlM88c, PJ85]. Checklist [Dav84]. Chemical [Gro89, Ulf85]. Childless [Bru85a]. children [Bru85a, Cam82, Kru84b]. China [BC80b, Che88, PJ85]. Chinese [Baa88, BC83b]. choice [Cro89a]. choices [Ahm80, For88b]. Chomsky [Gol85a]. chronological [Cam82]. Church [Bak83]. Churches [Hes82a]. Circular [BC81f]. cities [Col83]. citrus [Ano83d]. City [Col81a]. civil [AH86, Har87, Hes82b]. clarity [dR88]. classic [Mor84]. Classification [Boy86a]. climate [Man80b, Sta88]. Climatic [Dav87b]. cloth [O’T83, Syr83]. CMEA [Sin88c]. CMG [G.84]. Co [FGSA86, BM85a, Hry81, Iri85, Jon85a, Kap89, Kin87, Mar85, Nap88, Wat82a, Gil84, Gri84]. co-existence [Wat82a]. co-operation [Hry81, Kap89, Kin87, Mar85, Nap88]. co-operatives [BM85a, Iri85, Jon85a]. Co-ordinating [FGSA86]. Cogito [Pyk81]. cohesion [Sin87a]. Colemanic [San81]. Collaborative [Nap89]. Collapsing [Wat82a]. Collectives [Smi85a]. Colombia [Tal85a]. colony [Bul81]. Colourful [Pol89]. Combined [Jac87, Jen82b]. come [Dor86]. Coming [For86a, Gia86b, Smi85b]. commemoration [GM89b, Rav82]. commentary [Wil82]. Comments [Ano84b, Hes82a]. Commercial [Sim80, Ste82]. Commission [rpHK84, Mor80, Nar88]. Commissions [Vis85]. Committee [Pal87, Sin87a]. Commodity [Cam83]. Common [Hel83b, Der85, Tra83a]. Commonwealth [Gol86a, Hou82, Hou83, Kin87, Wil84b]. Communication [Gol84a, Wil84b, Fas84, Hil84b, Hil84c, Hor80, Mea86, PWG84]. Communications [Fit85a, Por84a, Mai89, Wil84a]. Communist [Che88]. Communities [rpHK84, FW82]. community [Irw89, CO80, Ste85]. companies [War85]. Comparative [GPH83, Sin88a, LLB86]. comparison [Dha84]. Compendium [BC83c]. competition [Mar85]. complaint [Gol85a]. comprehensive [Nel82, Pol89]. Computer [TC86, Kru84b, Tig83]. Computerization [Kir85]. Computers [Oak84]. con [Arn85]. concentration [Coh81]. concept [Hay84b, pAM81, Rea89]. conceptional [Pol82]. conceptual [GPH83, KM89e]. concern [KM89e]. conclusions [Pav84]. Conference [DB82, Ano80r, Emm80]. confidence [For88d]. confinements [Dav82]. conflicts [Arn83]. connect [Ver87b]. conscience [BC83f]. Conscious [Hay86a, Hay86b]. consensus [Der85]. consequences [Shi84, War89, Wat83b]. conservation [Gre82a, Pic80]. considered [Cam86]. Conspiracy [Hay89a]. Constraints [Col81a]. Consumer [Tsa89]. consumerism [BS83e]. contemplates [DeW82]. content [Swe86a]. Contentious [Gum82a]. context [Mul81, Sal81]. continent [Geo86]. Continuing [Ano85a, Ano86a]. Contradictions [Lic82]. Contribution [KM89b, Bul81, Las80]. contributions [Rot86]. control [Bel86, Gal87, Hug83, Tho82]. Controlling [Ash83, Jen82a]. controversies [Vis85]. controversy [BC81e]. conversion [Few80]. cooperation [Oak84]. coping [Gor84]. Copyright [Ano85b]. core [BS83b]. corporate [Irw85]. correction [Ano85b, Edg89]. corrective [Ash82]. Corrigenda [Ano81a]. Corrigendum [Ano83a]. costs [Bro83a, RF80]. countering [Lum82]. Counting [TS83a]. countries [Acc84, Ahm80, CO80, Ghe82, dMSG89, Gun80, Hil86a, LLB86, Mus80, ST85, Sin88c, Yoo86]. country [Bah85, Dem85, Gut80]. couples [Bru85a]. courts [BC82a]. Crash [Bro83d]. credit [Sim80]. crime [Sta83a]. crisis [Tal83]. Critical [Ano84c]. Critique [Gol81a]. Croom [Col83]. cross [Sta81]. cross-fertilizations [Sta81]. crossroads [Cox83]. crowded [BS84c]. Crowther [BC83g]. Cruise [Gum81a]. Culling [Sco89]. Cultural [Gol86b, Rah86]. culture [Ano84k, Fit85b, Rah89]. culture-bound [Ano84k]. culture-free [Ano84k]. cultures [Ahm86, Tor84]. cured [Tai81b]. cures [Smi85a]. Curing [San81]. Current [LJ83, Sør87]. Cushioned [Sta83a]. cuts [Lum82]. Cyprus [Ano83d].

D [Rav82, Ano84c, Bon84, Pic80]. damn [Swe86b, TS89]. Dangers [Dav82]. dark [Hil86b]. data [Mul87]. Dawn [Kel87a]. dealing [ZL87]. death [Sko86, Wat82b]. debate [BC81a, Hin81]. decades [For86c]. decision [BC80c]. decisionmaking [BS82b]. Decisions [BC82a, Gum81a]. decline [Bue80]. Defence [Rob81a, Bro83f]. Deference [JK82]. Defining [Fry82, Jos88]. definitions [Gol86c]. dehydration [Nag82]. Delhi [Tai83b]. delinquent [Bru85a]. Demand [For89]. demands [Col81b]. Democratic [Bra80]. Demographic [Hel81d]. Demythologizing [Gil83a]. department [Gre85]. deployment [Bri82]. depression [San81]. Derek [Gol83]. design [Kru84b, May88, Kru84a]. Designing [Chu87, Len88, Ros83]. determine [JS83]. determinism [Neu85b]. developed [Ahm80]. Developing [Ano81i, Dem85, Gia86c, Gun80, Acc84, Mus80, Yoo86]. Development [BC82a, DG88, Hoc86, Ano80i, Bow84a, Bro83b, Cam83, Cot86, Cro89b, Eis84, For85c, Ghe82, Gia80, Gol86b, Haq89b, Hel81d, Hil86b, Mac89a, NF82, OM87, Rah89, RM80, Smi85b, Tal83, Tho81, Wat82c, Wie81]. Dewitt [DeW82]. diagnosis [Dav87c]. dialectical [Csi82]. dialogue [Gia86a]. Diary [Ano87a, Ano88a, Ano88b, Ano88c]. diet [Bar84]. Dietary [Pyk87]. Difficult [Div88]. Dirge [Mun83]. dirigisme [Mun83]. Dirty [BS82a]. Disabled [Mel86, Kir85]. disarmers [GW83, Tra83b]. disarming [Ege89]. disaster [Coh81, PWG84]. Disconnecting [Por84b]. discovery [Sta89]. discrimination [Bak86]. Discussion [Ano81b, Ano81c, Ano81d, Ano81e, Ano81f]. Disease [Bur81]. diseases [BBGM85, LJ83, TS80, Vel81]. disillusion [Tol83]. disinterested [Bel86]. disorder [BM85b]. display [BS84c]. dispute [Sav85]. distinguished [Col82a]. distribution [Jac80]. divergent [Mac89b]. Diversity [Ehr80, Gum87b]. Divestiture [Por84b]. divine [Pyk80]. Division [Bon86a]. Do [BM85a, Cur81]. Dockyards [Wat83b]. doctrines [Pyk87]. Does [Med84]. doesn’t [Boy86b]. Doing [Duc83]. doors [Old89]. down [Hay87, Zim84]. Dr [Ano87f, Ash82, DeW82]. drift [pJAH83b]. drive [Boy83]. Drugs [TS83b, Sar80, Tai81b]. dumping [Sir85]. during [Kin87]. Dutch [BV81]. dynamic [Hag82]. Dynamics [Blu85, BC84].

early [TE89]. earthquake [Irw85]. East [Ano86b, Ano86c, Ano87b, Ano87c, Ano87d, Ano87e, Ano88d, Ano88e, Ano88f, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Moh86]. Eastern [Ano85c, Ano85d, Ano86d]. Ecocatastrophe [Hel81a]. ecological [NF82]. ecology [For85a, GPH83, Tho81, Wal83a]. Economic [Baj82, Cer80, Gia80, BC82a, Eki86b, Gol86b, Jas88, Jor86, Las80, Man80b, Nde85, OM87, Sal81, Sha88a]. Economics [Rah85a, Bud89b, Ray86]. Economies [Ano81i]. economy [Col88b, Csi82, For85a, Gia86b, Gia86c, Jas89, Sak86, dlM88c]. ecosystems [KM89e]. ecumene [Hou82]. ed [Col83]. edited [Ano84a, Hil84a, Por84b]. editor [Tra84]. Education [Duf84, Fas84, Ano80g, Ano84m, Cro84, Hor80, KM89b, Lal84, Sah80, Tol83, Wil84b, Yoo86]. Educational [Eis84, Kru84b]. EEC [Ano84e]. effect [Ids84, Mor81b]. effects [BC82b, ZL87]. efficient [Lan89]. Egypt [Hel81d]. eight [Pav84]. Eighteen [Hil86a]. electricity [Rüd85a]. Electronic [Hil84a, Cro84, Hil84c, Lal84, Shi84, Str84]. electronics [BM88a, Uen86]. Elements [Kru84a]. elephants [Sco89]. else [dlM88a]. embryos [Bar87b]. Emerging [Col81a, Boy85]. Emirates [BC82a]. emphasises [Bri82]. Empire [Hou82]. employment [Bon86b, Bro86, Duc88, Fai82, HB88, Jac80, Kir85, Pet84, SW86]. endangered [Sar80]. Energy [BC82a, Gil83b, GR88, Hes82a, Man80a, Pri83a, Pyk80, Swe87, Thr83a, BC81a, Bar87a, Bra80, Ell87, Fel81, KM89c, Low83, MS81, Pic80, Rah87, Sha84, Sør87, SS86, Thr83b, WW85]. engagement [dlM88d]. Engineering [Hel83a, Hry81, Thr81]. enigma [Zuc81]. enough [Gum81a]. Ensuring [KM89c]. enterprise [BS84b]. Environment [Hes82a, WM89, Bör86, BR86, Dro80, Mac89a, Sha89]. Environmental [Bow80, Hes82a, Hor80, Irw89, Sah80, War89, WW85, aH81, Ano80m, Bow84b, BC80c, BGM82, Nel82, RD80, Tal83]. environmentalists [BC82a]. equality [For89]. equity [Gil83b, Mor81b]. era [Gol88, Kel87a]. ergo [Pyk81]. escaping [BC81g]. establishment [BC84]. Estimation [Cam82, Fer82, Gri80]. etc [Gum87a]. Ethical [Hes82a]. Ethics [Pri85, Goo80]. Eugenics [Har89]. Eureka [Fel88, Nap88]. Europe [Ash83, Bro83e, Geo86, Gum87a, Här88a, HL88, Led85, Nar88, Pet84, Pon88, RD89, TE89]. European [DB82, Ano85c, Ano85d, Ano86d, Ano86b, Ano86c, Ano87b, Ano87c, Ano87d, Ano87e, Ano88d, Ano88e, Ano88f, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, CO80, Col86, For88b, For86b, Hes82a, Hoc86, rpHK84, Mac86, Nap88, Nar88, Pon88, Rub88, Ste85]. evaluated [Bud89a]. Evaluating [Mul87, Ram88b, Coz87]. Evaluation [SH86, Fer82, Nar87]. evaluations [Chu87]. Evaporation [Hug83]. Everyone [Hay89b]. evolution [Ano80q]. exaggerated [Bue80]. excellence [Ii82]. exchanges [RLH82]. Exemplary [Sho88]. existence [Wat82a]. exogenous [For85b]. Exotic [Vel81]. Expanded [Bah85]. expansion [Baj82, Eis84]. expenditure [AH86]. Experience [Ega88, Dem85, Iri85, Jas89, Rah87, RDU87, RCH87, Wri84]. Expert [Coz87]. Exploding [Hel81b]. Exploring [Gen84]. explosion [For86a]. export [Hes82b]. exports [Shi86b]. Extraordinary [dR88]. Extrapolating [BC81g]. extreme [JK81]. eye [Tit85]. eyes [Hil86b].

F [Ano88l]. fact [Lun87b, Swe86a]. factor [Duc83]. factors [Bri82]. Facts [Ano80e, Ano80f, Ano82a, Ano82b, Ano80d]. false [Seg88]. famine [Tan81]. far [Ano80c]. farm [Bal88, Haq89b]. farmer [Arn80]. farming [Dey84]. Fascinating [Zim89]. fault [dlM88a]. Fears [Bue80]. feasibility [Har87]. feed [Nag82]. fellah [Arn80]. Fells [BC81a]. Female [Fre81]. Feminist [Bak87b, Hay86c]. fertilizations [Sta81]. fiction [Hay86c, Lun87b]. field [Smi89]. fields [Boy86a]. Fifth [May87]. fifty [GM89a]. Figures [Ano80e, Ano80f, Ano82a, Ano82b, Ano80d, Ley88a]. Filling [Mla89, Wil83a]. Financial [Bha82, KM89e]. Finite [BC82a]. Finland [Ale85]. fire [dlM88e]. firm [Ram88a, Rot86]. firms [For85c, RD89]. First [DB82]. Fission [Hes82a, Sta89]. five [Ber89, Neu85a]. fixation [Pos87]. fluoridation [BGM82, BBGM85]. Fluorides [BGM82, BBGM85]. fluorocarbons [Lov81]. focus [Bar83b]. Following [Mor84]. folly [Nan82]. Food [Tan81, Ano84f, Gen84, Gil83c, Maz80]. forecasting [EMR^{+}80, Geo89]. foreign [Old89, Wei89]. Foresight [Geo89]. found [For88d]. Foundation [Ano84l, Ano87f]. Four [Neu85a, Yox87]. fragmented [Geo86]. frame [Tor82]. framework [Ano88g, KM89e, Rot85]. France [Har87, BR89]. Francophone [EDR85]. free [Ano84k, Wil89a]. Freeman [dlM88e]. French [Cha89]. fresh [Per85]. Friedenberg [Por84a]. Friendly [Kru84a]. frontiers [Gun80]. FRS [G.84]. fuel [RF80]. fuels [Few80, RCH87]. full [Duc88]. Functions [Rav82]. fundamental [Hay88b, Köö85]. funding [Tsa89]. Furniture [Syr83]. Further [Cam84]. futility [Wat83d]. Future [For88c, Rüd85a, Ano80r, BC82a, Car86, Fel88, Fit85a, Gol84b, Hod83, KM89c, Kin85, Mai89, Mar87a, McG89, Sør87, Hil84a].

G [Dav84]. game [Men80]. gap [Ein84, Fas84, For86d, Mla89]. gaps [Wil83a]. Gendron [Wil89b]. generation [Jac87, May87, Rüd85a, SW86]. genes [JK82]. genetic [BBGM85, Vis85]. genetically [Sha89]. gentle [(Re84]. germs [TS83a]. Getting [Mun81, Thr84]. Ghana [OM87]. Giant [Hay83]. glimpse [Mur86b]. global [Gre86, Tal83, Ano84a]. glowing [Bar83a]. goal [Ste85]. goals [Maz80, Tal83]. God [BC82a]. good [Van87]. goose [Wat83g]. Göteborg [DB82]. governance [Ale85]. Government [Ano88l, Cot86, Cro89a, Gum86a, BC80b, Gre85, Lei88]. Gower [Dav84, Hil84a]. gravity [Ano84b, Bra84]. Great [Bon86a, Köö85]. Greece [Dem85, Sin88b, Sot87]. greenhouse [Ids84]. Greening [Irw88a]. ground [Sir85]. group [MS81]. groups [WW85]. growing [Nag82]. Growth [Col81a, Nde85, Ano80b, Gre83, Kel82, Las80, Sta81]. guardian [Cri81]. Guide [Dav84, Fry80a, Mg81, RT84]. Guidedogs [BS82b]. Guidelines [Fal84]. Gulf [BC82a]. Gummett [Jon81a].

H [Mey82]. H-Bomb [Mey82]. Hamlet [Sin88c]. handicapped [GS85]. Handling [Hil84a]. happened [Gar87]. hard [Weh84]. harmony [KM89a]. Harold [Wal83b]. Harper [O’T83, Syr83]. Harry [Por84b]. Harvard [Ano88l]. hazards [Ulf85]. healing [(Re84]. Health [Eki86a, Sco89, TS86b, Bar84, Hel83a, RF80, Ulf85]. heat [Gri84, Jac87, Jen82b]. Heckel [Kru84a]. Helm [Col83]. help [Per85]. helping [Cam84]. Helsinki [Emm80]. hidden [Jor86]. hierarchical [Lam84]. High [Baa88, Col88a, Fit85b, Mal87, Seg88, Swe86a, Cha89, Cro89b, Jos88, Pow86]. high-temperature [Cro89b]. Higher [Ano80g]. Historians [Gol84b]. Historical [Pri86, Rah88]. History [Hay87, Rah80, Rah86, Wri84]. hoc [Ano80a]. holes [Mor81b]. home [Dav82]. homoeopathy [Man82]. hope [Tol83]. horizons [Bah85]. horticultural [Acc84]. Hot [BS83b, Bro83d]. housing [Per85]. Hoyle [Fry80a]. Human [Bar87b, Hil84c, PWG84, RM80, Tho81, Cah83, GPH83, Hil84b, KM89e, NF82, Smi87]. human-ecological [NF82]. humanity [WM89].

ideas [Ano85j, Has86]. Ideological [Div89a]. Ideologies [Lam84]. ideology [Wal83a, Wal83b]. II [Hes82a, Kin87, Rot81, Tra83a]. III [Por84b]. illusion [Gal87]. im [Hel83b]. images [BC83d]. imbroglio [BC80a]. immigrant [Tor84]. imminent [PWG84]. Impact [Kha87, Por84b, Bra81b, Fai82, Kir85, RD80, Shi86a]. impacts [Baj82, BC82a, Gol85b]. imperil [Cur81]. implications [BR86, Don84, Few80, Gre82a, Hel81d, Ids84, OM87, RD89, Wri84]. importance [Hay88b]. Imported [Shi86b]. inability [For85c]. Inaction [Nan82]. incidence [Dha84]. incident [BC83e]. incomes [Jor86]. Index [Ano80h, Ano81g, Ano82c, Ano84d, Ano86e, Ano87g, Ano88h, Ano89f]. India [Ahm85, BR86, BM88a, BM88b, Kha87, Rah85a, Rah86, Ram87]. Indian [BC83d, Rah85b, Rah85c, Rah87, Rao86]. Indicators [Eki86b, Ano84e, Fab84, Fal84, Gre85, Hil86b, Pav84]. indicted [Sha84]. indigenous [Haq89a, Haq89b]. Individual [Gol85b]. Indonesia [aH81]. Industrial [Mac89a, BR89, Hou87, NF82, Ray86, RD80, Rot86, dlM88c]. industrialised [dMSG89]. industrialization [Wad85]. industries [Cha89]. Industry [Lei88, Ano83d, BM88a, BM88b, Bro84, KM89f, Pal87, Rao86, Tig83, Uen86]. inefficiencies [Fer81]. infection [Gol84a]. influence [Gum86a, ZL87]. Information [Dav84, Hil84a, Smi87, BC81f, Ein84, Fai82, Len88, Oak84, Ram87, Sch84, Uen86]. initiatives [Bro86, For85b, SW86]. ink [Ash80]. innocence [Ege89]. innovate [BM85a]. Innovation [Bue80, Fry80b, Gre82a, Man80b, Sar80, Sim80, Wil82, Baa88, Duc83, Jas88, Jas89, Kod86, Lan89, Rot86, Stu88, Van88]. innovations [Gar87, Haq89a, Haq89b, RDU87, Wol85]. innovativeness [BM88a]. input [Bow84a]. Inquisition [Han84]. insights [Gum82b]. inspection [Cri81]. institute [Web89]. institutions [Bha82]. instruments [Ben87]. Intangible [Bud89a]. Integrated [FW82]. Integrating [HL88]. intelligence [Arn85, Jah86]. intelligent [Bak87a]. inter [Pow86]. inter-related [Pow86]. interacademy [RLH82]. Interaction [Baz84, Csi82]. Interactions [BC80b]. interdependency [Mar85]. interdisciplinarity [Mor81c]. interest [Pea81]. interface [Pig86]. intergovernmental [Emm80]. International [Ano80i, Shi86a, Bra81b, Bro83c, Fab84, Mai89, Sak86, Sin87a]. internationalism [Has86]. intervention [Dod89b]. Introduction [Ano84e, Hay84a, Gor84]. intuition [Mun82]. investigation [GPH83]. investment [Fry80b]. involvement [BV81]. Iran [Boy83]. Iranian [Tai84a]. IRDAC [RD89]. iron [Bro84]. Israel [RDU87, Tal80, Tal89, Tor84]. Israeli [Ste85]. Issue [Hes82a, G85c]. issues [BS83b, Gro89, Rub88, Sha88b]. Italian [Nap88].

J [Gol83, Hel83b, Hil84a]. J. [Rav82]. Jai [Rah85b, Rah85c]. Japan [BM88a, BM88b, Cam86, Cro89b, Cum87, Neu85a, Shi86a, Shi86b, SS86]. Japanese [BS83h, Dor86, Has86, Kod86, Mas86, Sak86, Uen86]. John [Ano84a, Ano88l, G.84]. Joseph [Ano87f, Gol87a]. Journalism [Ano84a]. Journey [Bai84]. Judith [Por84a]. justice [For89, Thr83a].

Kaplinksi [Sav85]. Kennedy [Ano88l]. Kenya [Eis84]. Kerala [EMR^{+}80]. kindled [dlM88e]. know [Bro83c]. know-how [Bro83c]. knowledge [KW87, Lin83, Tor82]. Korea [BM88a].

labor [Shi86a]. laboratory [DeW82, Sco89]. lagoons [Las80]. Land [Ano84f, Col81b, Wil83b]. Landon [O’T83]. Landscape [Hel83b]. language [Hay85]. languages [Cla88]. Lanka [Liy87, Ram88b]. Lankan [dS89]. lapdogs [BS82b]. Large [Dav87a, NF82, Rot86]. largest [Neu85a]. late [Gum81c]. Latin [Wat82c, Wat83e, Wat83f]. Law [Hel82a, Baz84, Mat83]. LDCs [Jac80]. Lead [Jon81b]. leaders [Dro80]. leadership [Mar85]. learning [Bah85]. leben [Hel83b]. Legal [BC82a, BS82a]. legislation [Che88]. legitimacy [Dod89b]. leisure [Col81b]. less [Ahm80]. less-developed [Ahm80]. Lessons [Cro89b, Kru84b, Liy87, Sco89, Rah80]. Letter [Tra84, Wil80]. level [BM88b, Ram88a]. liberty [Tam84]. libraries [Lum82]. library [Bro83a, Lum82]. Lies [Swe86b, TS89]. Life [BC82a, Cum87, Mar87a, Tam84, Sta86a]. Lifelong [Cro84]. Light [Gri84, Gum81c, Hil86b]. Like [Ver87a]. Limited [Mg81]. Limits [Col81a, Wei87, Ano80b]. Limps [BC82a]. line [NSA^{+}88]. Linking [Arn83, BR89]. Linköping [Ing87]. Links [Bar83a]. listings [Hay84b]. Literature [Ano87b, Ano85c, Ano85d, Ano85k, Ano86d, Ano86b, Ano86c, Ano86l, Ano87c, Ano87d, Ano87e, Ano87o, Ano88d, Ano88e, Ano88f, Ano88n, Ano88o, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e]. liver [Rob81b]. lives [Sho88]. living [Gre86]. Local [SW86, Bro86, Cha89, Irw88c]. Location [BS83d]. London [Col83, Syr83]. Long [Edg89, BS83a, Bue80, Hel81d]. long-term [Bue80, Hel81d]. Longmans [Ano84a]. look [Gum84a]. Lord [BC81b, Bul83, G.85b]. lore [Hel82a]. Lost [Tho85, Ros83]. Love [Mur86a]. Ltd [Bon86a]. ludicrosity [JK81]. lumbers [Wil80]. lump [Ver87a].

M [Por84b]. M. [BC83g]. Machrinhanish [G.85b]. Macroproblem [Arn84]. Mail [Hil84a]. Making [Har88b, Yoo86, Ahm85, BC80c]. Malaysia [BS83e]. Management [KM89d, Mar87b, WS83, Ano80j, Bow80, BS83d, BS83f, BS83h, Bru85b, Dav87c, Gri80, Lin83, Nel82, PJ85, aH81]. Management-union [Mar87b]. Manager [Dav84]. Manchester [Jon81a]. manipulated [Sha89]. manipulation [Vis85]. Mapping [Gar83]. Maps [San82]. market [Rea89]. Marshall [Syr83]. masculine [BM85b]. Mass [Ano84a, Wat82b, Don84, PWG84]. Master [Cha88]. material [Bar87b, BC81c]. materialism [pAM81]. materials [For88e, Pic80]. mathematics [Pow86, Wil84b, Yoo86]. matter [Arn89, Med84]. matters [Ano80m]. May [DB82]. MBE [G.85b]. me [Boy86b]. Meanwhile [Gum85]. Measurement [BM88b, Ram88a, Ley88b, Mul87]. measures [Bow84a, Jas88, Rao86]. Measuring [BM88a, NSA^{+}88]. mechatronics [Kod86]. Medal [Ano87f]. Medawar [O’T83]. Media [Ein84, Don84, PWG84, Ano84a]. Mediating [Bau87]. medical [Har89]. medicine [Ano80k, Coo83]. medium [RD89]. Meeting [Col88b, Emm80]. Meetings [Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano88i, Ano88j, Ano88k, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l]. megabucks [War87]. megawatts [War87]. Meiji [Ben87]. Mel [Syr83]. Merrill [Ano84a]. metacomputer [Gol84a]. method [Haq89a]. Methodical [pAM81]. meticulous [Irw88b]. Michael [Col83]. microcomputer [ST85]. Microcomputers [Hil84a]. microelectronics [Bra81b, Kha87]. Migraine [Tai81b]. Military [AH86, Ano80k, Bul83, Gal87, Gum87b, Har87]. mind [Arn89, Med84]. minds [Yox85a]. Miracle [BS84a]. misapprehensions [Edg89]. Misleading [San82]. miss [GW83, Tra83b]. missed [Wol85]. missions [Ege89]. mistake [Tra83a]. mobility [Sin87a]. model [Ahm80, Che88, Fel88]. modeling [JS83]. Models [Bro83c, GPH83]. Modern [Zim85, Ben87, Mur85, Pyk87]. modernization [Bal88]. Mondragon [Iri85]. money [Kec84]. monitoring [JS83, Lan89]. Moral [BC81c, BC83g, Yox85b]. motivating [Seg87]. much [BS83c, Gum81a]. MX [Gum81a]. My [Boy86b]. myth [BS84a, Jos89, Tho83].

National [Lic87, Mey82, Tor84, Ley88b, NSA^{+}88, Ram88b, Rot85]. nations [Gia86c]. Nato [Sin87a]. Natural [CO80, Ano80j, Bör86, KM89d, KM89f]. nature [Cah83, KM89a, Luc84]. Nazism [Har86]. near [Rea89]. necessary [pAM81]. need [BS83a, Coo84, Ii82, Ros83, Tal84]. Needham [Ano87f, Gol87a]. needs [Hor80, KM89e]. neglect [Jen82b]. neglected [For87, Lam84, Rah80]. neighbour [Mur86a]. Nepal [Hel81d].

Netherlands [Bro83b, Hag82, Vis85, vS88]. newly [dMSG89]. next [G.85a]. NHS [TS83c]. Nigeria [Ega88, Fam86]. Nitrogen [Pos87]. no [Hay88a, Hel82b]. Nobel [Dor86]. non [Ahm86, Bar87a, Haq89b, Sin88c, ZL87]. non-academic [ZL87]. non-farm [Haq89b]. non-Soviet [Sin88c]. non-western [Ahm86]. North [BS83f]. nostalgia [Seg88]. note [Ano88m]. Notes [Ano84g, Ano84h, Ano87m, Ano87n, Lun87a]. notion [Sta86a]. Nova [BC81e]. Nuclear [Bar83b, BC82a, Fel81, Hes82a, Mar86, Mar88, Pea81, Sha84, Sha85, Sim87, Bar87a, Blo83, Bra80, Bri82, BV81, Col84, Cur81, DeW82, Fry80a, Har87, Hin81, Hod83, Sta89, SS86, Tho85, War89]. numbers [Hel81b]. Nurturing [Ein84]. Nutrition [N’D80]. NY [Ano84a, Kru84a, Por84a, Por84b].

objectives [Pav87, Sch84]. Observations [Zuc81]. obsessions [Mel86]. Occidenting [BC83e]. Occupational [BR86]. occupations [Oke86]. OECD [Fab84, Rot81]. Office [Hil84a, Mat83, Smi87]. Official [Gum86b, Hes82a]. Oil [BC82a]. Old [Dic86, Boy87, Sch84]. On-line [NSA^{+}88]. Onchocerciasis [Tho82]. Only [Ver87b]. Onwards [BS83g]. Open [MS81]. opens [Old89]. operation [Hry81, Kap89, Kin87, Mar85, Nap88]. Operational [Sta81]. operatives [BM85a, Iri85, Jon85a]. Opportunity [Col81c, Kel87b]. Options [Bar87a, Ell87, JK81]. ordeal [Boy83]. order [Bro83c, Lam84]. ordinating [FGSA86]. Organise [Dav84]. organisms [Sha89]. organization [Hou87]. organizational [Ano83e]. Organizations [Str84]. Orient [Ben87]. orientals [BC83e]. orientation [Col81c]. oriented [Ahm80]. original [Dic86]. originality [Shi86b]. origins [Gol84d, Jos89, Wat83e]. our [Cur81, Rob83]. Out-of-school [Hoo86]. outcomes [Wat83e]. outlook [Cer80]. output [NSA^{+}88]. outsider [Jah86]. overdue [Edg89]. ownership [Der85].

P [O’T83]. Pacific [Hil86a]. Pacione [Col83]. pages [Col83]. pamphlets [Tai84b]. Panama [Tal85b]. Panel [Ano83e]. Paradigmatic [Wie81]. paradox [Cha88]. Parks [Lic87, Jos89]. parliament [Pon88]. Parliamentary [Pal87]. Part [vS88]. Part-time [vS88]. participation [Ano80m, Mar87b, Nap88, Nel82]. Partly [Ano85j]. Partners [Van88]. partnership [Coo83]. past [Fit85a, Mur86b]. Patent [rpHK84, Ste84]. Paternalistic [Smi86]. path [(Re84]. paths [BC82a]. patronising [Moh86]. pattern [Mag84, Mor84]. Patterns [Ano80l]. Paul [Kru84a, Ano84b]. PC [G.85b]. Peace [Moo89, Rob83, Bul81, Bul83]. peacetime [Gum89]. Pen [Ash80]. Pen-and-ink [Ash80]. people [Irw88c, Che88]. perception [Bul83]. Performance [Gre85, Ley88a, Ley88b, SH86]. Pergamon [Por84b]. periphery [Fro84, SH86]. personal [Duc83, Rog81]. perspective [Col86, Gil84, GC87, dMSG89, Jah86, Moo86, NF82, Pri86, Rah88, Sak86, dS89]. perspectives [Ram87]. Peru [Tal84]. petroleum [RCH87]. Philip [Jon81a]. philosopher [Div89b]. philosophy [BC82b, Irw86, N’D80]. physical [Ano80b]. Physicist [Div89b]. physics [Mag84]. pieces [BC81d]. place [Kem87a, Kem87b]. planned [Jas89, Wat83d]. Planning [Ano81i, Col83, Hel81d, Liy87, Sah80, dS89, Hil84a, Hil84a]. plant [Web89]. Playing [Arn88a]. plea [Bar84]. plutonium [Hes82b]. Poems [Ano81h]. point [Arn88b, GW83, Tra83b]. Poland [Jas88, Jon85a, KW87]. Policies [Bon84, BC82a, Gor84, BC81c, Csi82, LLB86, May87, Sin88a]. Policy [Ano82g, Ano84l, Ano88l, Bul83, Hes82a, Yox85b, Ahm80, Ahm85, Ano80q, Ano84e, Ano88g, Bar84, Bar87a, Bha82, BR86, BS83j, Bro83b, Car86, Cha89, Che88, Cro89a, Dav82, Dav88, Dod89b, DG88, Fas84, Fre84, G.85a, GR88, Gol80, dMSG89, Gum81b, Gum86a, Hel81d, Hoc86, Ids84, Ii82, Irw89, JS83, Jas88, Jas89, Jon85b, Lan89, Lov81, Mul81, Pav87, Rot85, RD89, Sal81, Sha88a, Sha88b, Sot87, SS86, Tal80, Tal84, Tal89, Thr83b, Wil80, Wil83b, Wil89a, Wil89b, dlM88b, dlM88e]. policy-oriented [Ahm80]. political [Hou82, Lei88, Mor81c, Por84a]. politician [Fre84]. Politics [Ahm85, Bro83f, Csi82, For85a, Jos88, Lun87a, Mar88, Mep89]. Pollution [Jon81b, Bör86, JS83]. Poor [Kel84]. popularization [Bis80]. Population [BS83j, Ano84f, For86a, Gil83c, Gre83]. populations [Nag82]. portrait [Ash80]. Portugal [GC87, Sin88b]. Positivist [Mor81c]. posset [Hel82b]. Possibilities [Lal84]. possible [BC82b]. Post [Mat83, Coh81, Hou87, Sab80]. post-concentration [Coh81]. post-industrial [Hou87]. post-UNCSTD [Sab80]. potato [Nag82]. potential [Jon85a, Jor86, Lak81]. poverty [For86d]. Power [BC82a, Hes82a, Baj82, Ell87, Fry80a, Gal87, Hod83, Hou87, Jac87, Jen82b, Pea81, Sha85, Tho85]. pp [Ano84a, O’T83]. Practical [Seg87]. practice [Iri85]. practices [Shi86a, Por84a]. Praeger [Por84a]. Pre [Pol82]. Pre-conceptional [Pol82]. precipitation [Ash82]. preferences [Tsa89]. prehistory [Ulf85]. prescription [Ano80l, Hel82b]. present [Fit85a, Ulf85]. presents [Ash82]. presidential [Dav87a]. Press [Jon81a]. pressure [Gre87]. prevention [Bur81]. Preventive [Hel82b]. Price [Gol83]. prices [Cam83]. Primary [Rob81b]. Prince [Sin88c]. Principles [Por84a]. priorities [Bal88, Tsa89, WW85]. private [Jac87].

Privatisation [Web89]. Prizes [Dor86]. pro [Arn85]. Problem [Hay88a, Ash82, Fry82]. problematic [Gre82a]. Problems [Hes82a, Ley88b, Lin83, Pet84, Sha88b, Yap83, Ano85j, BC80c, LJ83, Oke86, RM80, Tit85, Col83]. process [Jas88, Nde85, Yox85b]. Processing [Hil84a]. Product [Sta86a]. production [Acc84]. productive [Gia86c]. productivity [Gol86c]. profession [Har89]. professors [vS88]. profits [Wol85]. Program [Ano88l, Baj82, Chu87]. programme [Oak84, Sin87a]. programmes [Har87, Ram87]. Programs [Fre81, Coz87, Kru84b, Mul87, Nar87]. Progress [Coo84, Mac86, Cri81, Eki86b]. Progressive [Mey82]. Project [Ano81i, Dav87c]. Projects [FW82, Dav87a, Fal84, Nap89]. promise [For88c, Tho85]. promises [Wat83f]. promoting [Haq89a, Sin87a]. Promotion [Haq89b]. Promotional [Rao86]. proofs [Pet84, Yap83]. Prospects [Blo83, Wat82c, Kel84, Oke86]. protect [Cur81]. Protectionism [Tig83]. proved [Kib81]. Providing [Pow86]. Proving [Arn88b]. psychical [Has84]. psychological [Bri82]. Public [Ano80m, Ano88l, Ber80, Hes82a, Jas88, KW87, Nel82, Ano84c, Bon80, Boy86b, Gum80, Hel81d, Ids84, Jas89, Lan89, Pea81, RD89, Ste82, Web89]. Publications [Ano80n, Ano80o, Ano80p, Ano81j, Ano81k, Ano81l, Ano81m, Ano82d, Ano82e, Ano82f, Ano83b, Ano83c, Ano84i, Ano84j, Ano84n]. purity [Mor86]. Putting [Kin85]. puzzles [BC83b].

Qualified [TS81a]. Quality [Smi88, BGM82, Pow86]. quantity [Smi88]. Quebec [Dav88]. Quest [Gol81b, Csi82]. Question [BC81b, Gri84, N’D80]. Questions [Hil81, ZL87].

R [Ash82, BC83g, Ano84c, Bon84, Pic80]. R&D [Bon86a, AH86, Cum87, GR88, Gre87, HL88, Lei88, Sta86b, Tal85a, Tal85b]. race [Col84, Cro89b, Dav87a, DeW82, Wat83c]. Radical [GD84]. radioactivity [CO80]. Rain [DB82]. Rare [Ray86]. rational [Csi82]. Rationale [Wri84]. rationality [KM89e]. Rationalized [Mun82]. raw [Pic80]. raw-materials [Pic80]. re [Bro83g]. re-appraisal [Bro83g]. Readable [Dod89a]. Reading [BS84c, Gum82a]. realist [Gil83c]. Realistic [Maz80]. reality [Bul83, Tho83, Tol84]. reasoning [Yox85b]. receive [Sim80]. recipients [Tai83c]. recognisable [Bud89b]. reconstructed [BC83b]. records [RT84]. recovery [Ano83d]. reductions [Blo83]. Redundancy [pAM83, Wor89]. Reentry [Fre81]. Refinish [Syr83]. Reflections [Ada82, Hin81, Sin87b]. reform [Jas88]. refraction [Mor81b]. region [Gor84, Hil86a]. Regional [Mor85, Emm80]. regions [RDU87]. Regulation [Sha89, Wie81]. Reindustrialization [Rot85]. related [Pow86]. release [Sha89]. relevance [Ben87]. relevant [Jas89, Yoo86]. reliance [Gol80]. Religion [Coo83]. Remythologizing [Har82]. Renaissance [Rah85b, Rah85c]. Renewable [Ell87, Rah87]. reorganisation [Cot86]. Repair [Syr83]. report [Rot81, Wil89b]. Reports [Hes82a, Tra83a]. Republic [Che88]. required [Cla88, Man80b]. requirements [Col88b, Sør87]. Research [Bow84a, Dod89b, LLB86, Ale85, Bal88, Bar87b, BR89, Chu87, Cot86, Cox83, Csi82, Dem85, DST87, Fam86, FGSA86, Has84, Irw88a, Low83, Lum82, MS81, Mul87, Nar87, Oak84, Pos87, Rea89, Sin87b, Sta88, Sta81, Tit85, Tsa89, Web89, ZL87]. resignation [Stu88]. resource [Nel82, RM80]. resources [Ano80j, Ano84m, Bör86, BS83j, KM89d, KM89f, Lin83, Wat83b]. respect [Sta88]. response [Eng80, Tal83, TS86b]. responsibilité [Don80]. responsibilities [Ber80]. responsibility [Sta89]. results [RD89]. Rethinking [ST85, SW86]. Reupholster [Syr83]. Review [Ano84a, Col83, Dav84, Hil84a, Jon81a, Kru84a, O’T83, Por84a, Por84b, Syr83, Coz87]. reviews [Ano84c]. revisited [Wil84a]. revolution [Cro84, Hil84c, Lal84, Shi84, Wil84a, Tai84a]. rice [Dey84]. rise [Col88a, War85]. rises [BC81a]. rising [BS83h]. Risk [Eng80, Fer82, BC80a, BS83f, GB80, Gri80, Gro89, Gut80, Irw87, JS83]. risks [Gia80, Goo80, Sut80]. rite [Mep89]. river [Sah80]. river-basin [Sah80]. Robert [Por84a]. Role [Bro86, Acc84, Cri81, Jon85a, Zim85]. Roots [Tai84a]. Row [O’T83, Syr83]. Royal [Wat83b]. ruled [Rüd85b]. rulers [Rüd85b]. Rules [Men80]. rural [FW82, Haq89b, NF82].

S [Por84a, Fab84, Gol84c, Kel82]. S&T [Che88, Hil86b, Sha88b]. safe [Hil85, Kem87a, Kem87b]. safety [Cri81, Zuc81]. sandwich [For86b]. Saving [War87]. Say [Fit85c]. scale [Jac87, NF82]. Scandinavia [Ash82]. Schizophrenia [Tho83]. scholars [Ben87]. School [Ano88l, Hoo86]. Science [Ada82, Ano82g, Ano84k, Blu85, BM85b, Dav88, For87, Ghe82, Gol82, GC87, Gow81, Hag82, Ids84, Irw84, Lan84, Led85, Lic87, Lun87b, Mep89, Mul81, Pal87, Pav84, dSP89, RT84, Rah86, Rah88, Rah89, Ram87, Rav82, Sal81, Sin87a, Sin88c, Sin88b, Sot87, Tal80, Tal89, Wad85, Wil80, ZL87, Zim84, dS89, Ahm85, Ahm86, Ano80q, Ano80r, Ano84e, Ano85a, Ano86a, Ano88g, Bak86, Bau87, Bel82, Ben87, Bis80, BC80b, BC81h, BC82b, Bro83b, Cha89, Chu87, Csi82, Cum87, DG88, Eis84, EDR85, Fal84, For86c, Fre84, G.85a, Gar83, Gol85b, Gol86a, Gol87b, Gol88, Gre85, Gum81b, Gun80, Har88b, Hay86c, Hoc86, Hoo86, Ii82, Jon85b, Kin85, Köö85, Lan89, Ley88a, Liy87, Mac86, Mar88, May87, Mea86, pAM81, Oke86, PJ85, Phi88, Pig86]. science [Pow86, Pri85, Ram88b, Rub88, Sin88a, Sta81, Sta83b, Ste85, Tal84, Tho81, Tor84, Wei87, Wei89, Wil84b, Wil89a, Wil89b, Wor89, Yoo86, Zim85, dlM88b, Ano84l, Ano88l]. science-based [Oke86]. science-society [Wor89]. Sciences [Hes82a, Baa88]. Scientific [Ben87, Mor81a, Ano83e, Boy86a, Col88b, Don80, Gol84d, Kin87, KW87, Ley88b, NSA^{+}88, Sho88, Sin87a, Sin87b, Sta88, SH86, Thr83b, Tit85, Tsa89]. scientist [Ber80, Sta89, O’T83]. Scientists [Bon80, BS82b, Fre81, Gum80, For87, Old89, Tor84, Jon81a]. Scientology [Hay88b]. Scintillating [Bon81]. Scotia [BC81e]. Scottish [TS86a]. Search [Gei86]. Second [Gum84b]. secrecy [Sim87]. Secret [Mey82]. Secretaries [Dav84]. secrets [Gum86b]. sector [Haq89b, Web89]. Security [BC82a, Mey82, Rob81a]. seeing [Hil86b]. selection [Pol82]. selective [Har86]. self [Bru85b, Gol80, Per85]. self-help [Per85]. self-reliance [Gol80]. seminar [Ano84e]. Sensationally [TS84]. Separating [Har87]. serious [Mor85]. servant [Gal87]. service [Bon80, Gia86b, Gia86c, Gum80]. Setting [BS83h]. Severed [BC81d]. sex [Pol82]. Sexual [Bak86]. sexually [LJ83]. shameful [For85b]. Shaping [Gre83]. Shipping [Sut80]. Shop [ZL87]. should [Sim80]. side [Gum82b]. significance [Emm80]. silence [Hay89a]. Silicon [Jos89]. Simpson [Hil84a]. since [BC80b]. Singapore [Hoo86]. Singh [Rah85b, Rah85c]. Sir [G.84]. Sitz [Hel83b]. six [LLB86]. sized [RD89]. skill [Ros83]. slanted [Pri83a]. Slavery [Gol84c]. slaves [Thr83a]. Small [Rot86, Jac87, RD89]. small-scale [Jac87]. sobriety [TS86a]. Social [Gol82, Hes82a, Joh86, Rav82, Shi84, Baj82, BC82a, Goo80, Ing87, Mor81b, Rah88, Wol85, Zim85]. socialism [Tai83a].

Socialist [Ano84n, Rub88, TS80]. Society [BC82a, Bau87, Hou87, Irw84, Len88, Mas86, Pig86, Sta86b, Uen86, Wor89]. socio [OM87, Sal81]. socio-economic [OM87, Sal81]. Sociology [Sta81]. soft [Weh84]. Software [Kru84a]. Solla [Gol83]. solution [Hay88a]. Some [Ano83e, Don84, Gar87, Gre82a, Hel81d, JK81, Ram87]. someone [dlM88a]. Sooshan [Por84b]. sophistication [Gol84e]. soul [Wat82b]. Sound [Fre88]. Source [Wat83g]. Sources [Dav84, Mul87, Rah87]. Soviet [Sin87b, Sin88c, Tor84]. space [Gow81]. Spaces [Col82b]. Special [G85c]. species [Sar80]. speculations [Hil81]. spending [Lum82]. Sphinx [Tai81a]. spiritual [Rah88]. sponsorship [Mag84]. SPP [Ano88h, Ano89f]. spread [TS80]. spruce [BC81e]. squishy [Weh84]. Sri [Liy87, Ram88b, dS89]. staff [Pow86]. staffing [Col88b]. Staggering [Ege89]. standards [Gre86]. Standing [Neu86]. stars [Tam84]. state [Ehr80, For88a, Gum81b]. Statecraft [BC83a]. Statements [Hes82a]. States [BC82a, Bru85b, Cro89b, Hel81c]. statesman [Div89b]. statistical [TC86]. statistics [rpHK84, Ste84, Swe86b, TS89]. status [Sør87]. steady [Ehr80]. steel [BM88b]. steps [Hay83]. Stewart [Bri82]. stigma [Wol85]. storage [Nag82]. Strategic [Jon85b, Wat83a, FGSA86]. strategies [Cro89a, Hil86b, LLB86, Sha88b, WS83]. strategy [Ano80i, Hou83, JS83, Kin85, Phi88]. Stratospheric [Gre82b]. stream [Thr84]. strife [Gre82b]. Striking [Fer81]. structural [Tal84]. struggles [Div89a]. STSA [Lak81]. student [BFSL85, Yox85a]. studies [Chu87]. study [Bis80, BM88b, BS83e, EMR^{+}80, Har87, Ram88b, Sho88, Sta81, Tor84, Wal83a]. Studying [DST87]. stuff [Dic86]. success [TS81a]. suicide [Arn88a]. sum [Pyk81]. sun [BS83h, Pri83b]. super [Bak87a]. super-intelligent [Bak87a]. superconductivity [Cro89b, Mal87]. Supernatural [Sta83b]. Supply [Gum82b, Gil83c, Sør87]. Supply-side [Gum82b]. Support [Lak81, Lei88, Pow86]. suppression [Mar86]. Surprising [Boy88]. Survey [Ano84a, BFSL85, Led85, TC86]. survival [Bel82, Gow81, Lun87a]. Surviving [Mor80, Nix87]. sustainable [BC82a, KM89c, Mac89a, Sta86b]. sustained [Haq89b]. Svedenstierna [Bro84]. Sweden [DB82]. sympathy [War86]. symposium [Sab80]. syndrome [Coh81]. Synthesis [TS81c]. system [Baj82, Fab84]. Systematic [Wal83b, Wal83a]. systems [Dey84, Ros83, RF80, Smi87].

T [Fab84, Gol84c, Kel82]. tale [NSA^{+}88]. task [Tol83]. Tax [Bra81a]. Taylorisation [Smi87]. Tbilisi [Emm80]. teachers [Gol86a, Pow86]. teaching [Lum82, Pow86]. tecbnology [Smi87]. tech [Cha89, Seg88]. Technical [Ano88m, Jac80, KM89e, Wol85, Col88b, WS83]. Techniques [Dav84, Nar87]. techno [Sha88a]. techno-economic [Sha88a]. technocrat [Ash83]. Technological [Ahm80, Cri81, Irw87, RDU87, TE89, For88b, Jen82a, Mor81a, Nap88, Nar88, Nde85, Neu85b, Pon88, Sco89]. technologies [Ein84, Gor84, Sch84]. Technology [Ano88l, Bon86b, BC83f, BS83a, Bud89b, Col84, Cro89b, FW82, GS85, dMSG89, HB88, Hes82a, Ing87, Joh86, KM89f, Mar85, Mor81b, OM87, Pet84, RD89, Van87, dlM88e, Ahm86, Ano80r, Ano84e, Ano88g, Baa88, Baz84, Bha82, Bis80, BM88b, BFSL85, BS83d, BS83i, Bro83c, Bro83f, Car86, Col86, Coo84, Cro89a, Csi82, Dav88, Dod89b, DG88, Fai82, Fal84, Fit85b, For85a, For86c, G.85a, Ghe82, GC87, Gre86, Gum87b, Gun80, Hil86a, Hoo86, Jon85b, Jos88, Jos89, KM89d, Kod86, Lan89, Led85, Lei88, Liy87, Mac86, McG89, Mur85, Oak84, Pav84, Pav87, Pig86, Pow86, Pri85, RT84, Rah88, Ram88b, Ram88a, Ram87, Rao86, Rot85, Rub88, ST85, Shi86b, Sin88b, Sør87, Ste85, Tal80, Tal89, Tho81, Wad85, Wei89, Wil84b]. technology [Wil89a, Yoo86, dlM88b]. Technology-based [RD89]. Teilhard [Hou82]. Telecom [Mat83]. telecommunications [McG89, Hil84a]. television [Ste82]. temperature [Cro89b, Mal87]. ten [ZL87]. tension [Gal87]. term [Bue80, Hel81d]. terms [For86a, Smi85b]. terrorism [For88a]. test [Jac87, Sim87]. test-case [Jac87]. Text [Boy85]. Thailand [TC86]. their [RD89]. them [Gar87]. theme [Dic86]. theological [JK81]. theory [Bow84b, Gia80, Lam84]. there [Ano80b]. thesis [TS80]. things [BS83c, Tai81a]. thinks [Tai81a]. Third [Col83, TE89, BM85a, Bow80, Cam84, For85a, Gar83, Hug83, Pen80, Per85, ST85, Sha85, Sir85]. thought [Mor85]. Thoughtful [Irw88b]. Thoughts [Bul83, Cam84]. threat [Köö85]. Three [Tai84b]. Thring [Tai81a]. thurowly [Cam86]. Tibet [Old89]. tidal [Ell87]. time [Bai84, Gum84b, Rob83, vS88]. Titans [BS83i]. together [Duc83, Mun81]. Too [Gum81c, Ano80c]. Totally [dR87]. Tracing [Gol84d]. trade [Wil89a]. Tradition [Mus80]. transfer [Bro83c, Hil86a, Lin83, Mur85, Nde85, ST85, TE89]. Transformation [Tor82, Nde85, Rah86]. transition [Bru85b]. transmitted [LJ83]. transnational [For85c, War85]. Transplantation [EDR85]. transport [RCH87]. trap [Pri83b]. trees [Col82a]. trendiness [pJAH83a]. Trends [pJAH83a, Ano83e]. Trent [Por84a]. Tribute [Gol87a]. Trident [Gum81a]. Triumph [Arn89]. trust [Sta88]. Truth [Gei86]. Turkey [Sin88b]. turned [Hay87]. TV [Kru84b]. Twentieth [Gol84e, Han84]. Twenty [Gol87c, Ber89]. twenty-five [Ber89]. Two [For86c].

UK

[Dav84, Ano80g, Ano80q, Ano82g, Ano85a, Ano85k, Ano86a, Ano86l, Ano87o, Ano88n, Ano88o, Bar87a, Cot86, Cox83, Gum81b, Har87, Hil84a, Mat83, Phi88, Rea89]. Ukraine [DST87]. Uncertain [BC80c]. uncertainty [Gia86a]. unconvincing [dR87]. UNCSTD [Sab80]. understand [Boy86b]. understanding [PJ85]. unemployment [Duc88, Rog81]. Unesco [Ano83e, May88]. Unfinished [War86]. unhappy [Tai83c]. Unhealthy [TS81b]. Union [Sin87b, Mar87b]. unique [BS83e]. United [BC82a, Bru85b, Cro89b, Hel81c]. University [Jon81a, BR89, Lal84, ZL87, Ano88l, Lum82, MS81]. unjustified [Jen82b]. Unlimited [Bar83c]. unresolved [Lic82]. Untaming [Luc84]. upside [Hay87]. upwards [BS83g]. Urban [Col81a]. Urbanization [San82]. USA [Ano84a, Kru84a, Por84a, Por84b, Bue80, Man80b, RLH82, Sar80, Sim80]. usable [Sap89]. usage [TC86]. Use [Bör86, Col81b, For85c, Wil83b]. useful [Sta81]. USSR [RLH82]. utilization [DST87, Sta86a]. Utopia [Gen84].

V [Por84a]. Valley [Jos89]. Value

[Bra81a, BC83g]. Values

[BC81h, Thr81, Bar83c, Bro83g, Car86, Gol86b, Rah88, Tal83]. variable [Sta86a]. Varied [Mac89b]. vehemence [Bak87b]. Venezuela [Bal88]. very [BS83c]. view [Ash82, BS83a, Gil83c, Mur85, Nar88, Pon88, Rog81, Tit85, Wil89b]. views [Mac89b, Pri83a]. vision [Rav82, dR88]. Visual [Kru84b]. Voices [Irw88c]. Volta [Tho82]. Volume [Ano85l, Hes82a]. votes [Bra80]. waiting [For85b]. Walsby [Wal83b]. want [BC83h]. War [BC82a, BC83h, Kin87, Bar83b, Bon81, BC83f, War89]. Warner [Kru84a]. wash [pJAH83b]. waste [Wat83b]. Wasted [Kel87b]. water [BGM82, BBGM85, Lin83, aH81]. Wave [Neu85b]. way [Fel81, For85b]. wealth [Eki86a, Gia86c, Gol86c, TS86b]. Weapons [Gum89, Rüd85b, Bri82, Cur81, DeW82, Hin81]. welcome [Boy83]. West [Moh86]. western [Ahm86, Ahm86, Bro83g, Has86]. westernization [Boy83]. Where [Dor86, GW83, Tra83b]. Which [For85b]. white [Sco89, Sko86]. Whitehall [Jon81a]. Who [Low83, TS81b]. wide [Kap89]. Widening [For86d, Man82]. widespread [Few80]. will [BS83c, Dor86]. win [Hay89b, Hou83]. win-win [Hou83]. windships [Nan82]. winter [Mar88]. wisdom [Nan82]. within [DeW82]. without [BC83e, BS84a, Sin88c, Tai81b]. Woman [Cah83]. Women [Dey84, Oke86, Acc84]. Woods [Col82a]. Word [Hil84a]. Work [Tol84, Dro80, Kin85, MS81, (Re84, Wal83b]. worker [BM85a, Bru85b, Jon85a]. workplaces [Ulf85]. workshop [Emm80]. World [Ano84a, BC82a, Col83, Gil83c, Gol80, Kap89, TE89, Ano80r, Ano84n, Arn88a, Gre82a, Gum84a, Kem87a, Kem87b, Mas86, BM85a, Bow80, Cam84, Ega88, For85a, Gar83, Hug83, Kin87, Pen80, Per85, ST85, Sha85, Sir85]. World-wide [Kap89]. worldview [Spe81]. write [Fit85c]. wrong [TS84].

year [ZL87]. years [Ber89, Bon81, Gol87c, GM89a]. York [Ano84a, Kru84a, Por84a, Por84b]. Young [O’T83]. youth [Yoo86].

Zuckerman [BC81b].

References

Accati:1984:WRH

[Acc84] Elena Garibaldi Accati. Women’s role in horticultural production in developing countries. Science and Public Policy, 11(4):219– 226, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Adams:1982:RBS

[Ada82] Sir John Adams. Reflections on a big science. Science and Public Policy, 9(2):81–87, April 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

al-Hajj:1981:EAW

[aH81] Mardjono Notodihardjo al Hajj. Environmental aspects of water management in Indonesia. Science and Public Policy, 8(3):181– 184, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Acland-Hood:1986:MCR

[AH86] Mary Acland-Hood. Military and civil R&D expenditure. Science and Public Policy, 13(1):52–54, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ahmad:1980:TCL

[Ahm80] Rashid S. Ahmad. Technological choices in less-developed countries: a policy-oriented model. Science and Public Policy, 7(2): 120–123, April 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ahmad:1985:PSP

[Ahm85] Aqueil Ahmad. Politics of science policy making in India. Science and Public Policy, 12(5):234– 240, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ahmad:1986:WST

[Ahm86] Aqueil Ahmad. Western science and technology in non-western cultures. Science and Public Policy, 13(2):101–105, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Alestalo:1985:CGR

[Ale85] Marja Alestalo. Changing governance and research in Finland. Science and Public Policy, 12(5): 279–285, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:AH

[Ano80a] Anonymous. ad hoc. Science and Public Policy, 7(1):74– 79, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:TPL

[Ano80b] Anonymous. Are there physical limits to growth? Science and Public Policy, 7(1):35– 42, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:BTF

[Ano80c] Anonymous. A bridge too far. Science and Public Policy, 7(4):302– 303, August 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:FFa

[Ano80d] Anonymous. facts in figures. Science and Public Policy, 7(1):70– 74, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:FFb

[Ano80e] Anonymous. Facts in figures. Science and Public Policy, 7(2): 152, April 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:FFc

[Ano80f] Anonymous. Facts in figures. Science and Public Policy, 7(3):229– 230, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:HEU

[Ano80g] Anonymous. Higher education in the UK in the 1990s. Science and Public Policy, 7(1):54– 61, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:I

[Ano80h] Anonymous. Index 1980. Science and Public Policy, 7(6):486–488, December 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:IDS

[Ano80i] Anonymous. International development strategy for the 1980s.

Science and Public Policy, 7(3): 214–216, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:MNR

[Ano80j] Anonymous. The management of natural resources. Science and Public Policy, 7(1):43–49, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:MM

[Ano80k] Anonymous. Military medicine.

Science and Public Policy, 7(4): 301, August 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:PP

[Ano80l] Anonymous. Patterns of prescription. Science and Public Policy, 7 (3):218–219, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:PPE

[Ano80m] Anonymous. Public participation and environmental matters. Science and Public Policy, 7(6):448– 469, December 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:Pa

[Ano80n] Anonymous. Publications. Science and Public Policy, 7(1):69, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:Pb

[Ano80o] Anonymous. Publications. Science and Public Policy, 7(3):223– 229, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:Pc

[Ano80p] Anonymous. Publications. Science and Public Policy, 7(5):398– 399, October 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:USB

[Ano80q] Anonymous. The UK science board: evolution of policy 1969– 79. Science and Public Policy, 7 (1):49–54, February 1980. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1980:WCS

[Ano80r] Anonymous. The world conference on science and technology and the future. Science and Public Policy, 7(1):61–65, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:C

[Ano81a] Anonymous. Corrigenda. Science and Public Policy, 8(3): 256, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:Da

[Ano81b] Anonymous. Discussion. Science and Public Policy, 8(2):91– 98, April 1981. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:Db

[Ano81c] Anonymous. Discussion. Science and Public Policy, 8(2):99– 109, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:Dc

[Ano81d] Anonymous. Discussion. Science and Public Policy, 8(2):128– 135, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:Dd

[Ano81e] Anonymous. Discussion. Science and Public Policy, 8(2):143– 147, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:De

[Ano81f] Anonymous. Discussion. Science and Public Policy, 8(2):151– 154, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:I

[Ano81g] Anonymous. Index 1981. Science and Public Policy, 8(1):ii– v, February 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:Pb

[Ano81h] Anonymous. Poems. Science and Public Policy, 8(1):80, February 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:PPD

[Ano81i] Anonymous. Project planning for developing economies. Science and Public Policy, 8(3): 251, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:Pa

[Ano81j] Anonymous. Publications. Science and Public Policy, 8(1):77– 79, February 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:Pc

[Ano81k] Anonymous. Publications. Science and Public Policy, 8(2):166– 167, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:Pd

[Ano81l] Anonymous. Publications. Science and Public Policy, 8(3):254– 256, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1981:Pe

[Ano81m] Anonymous. Publications. Science and Public Policy, 8(5):419– 420, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1982:FFa

[Ano82a] Anonymous. Facts from figures. Science and Public Policy, 9(1):

51–53, February 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1982:FFb

[Ano82b] Anonymous. Facts from figures.

Science and Public Policy, 9(3): 163–167, June 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1982:I

[Ano82c] Anonymous. Index. Science and Public Policy, 9(6):335–336, December 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1982:Pa

[Ano82d] Anonymous. Publications. Science and Public Policy, 9(1):48– 50, February 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1982:Pb

[Ano82e] Anonymous. Publications. Science and Public Policy, 9(2):111– 112, April 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1982:Pc

[Ano82f] Anonymous. Publications. Science and Public Policy, 9(5):280– 281, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1982:USP

[Ano82g] Anonymous. UK science policy bibliography. Science and Public Policy, 9(4):224, August 1982.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1983:C

[Ano83a] Anonymous. Corrigendum. Science and Public Policy, 10(5): 264, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1983:Pa

[Ano83b] Anonymous. Publications. Science and Public Policy, 10(2):97– 102, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1983:Pb

[Ano83c] Anonymous. Publications. Science and Public Policy, 10(5):265– 270, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1983:RCI

[Ano83d] Anonymous. The recovery of the citrus industry of Cyprus. Science and Public Policy, 10(6):301–303, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1983:SST

[Ano83e] Anonymous. Some scientific trends, some organizational aspirations: a Unesco Advisory Panel. Science and Public Policy, 10(2): 71–76, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:GJS

[Ano84a] Anonymous. Book review: Global Journalism: a Survey of the World’s Mass Media edited by John C. Merrill Longmans, New York, NY, USA, 374 pp, £17.50. Science and Public Policy, 11(6): 395, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:CPB

[Ano84b] Anonymous. Comments on Paul Brazier’s anti-gravity. Science and Public Policy, 11(5):279–280, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:CRP

[Ano84c] Anonymous. Critical reviews of public R and D. Science and Public Policy, 11(4):183, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:I

[Ano84d] Anonymous. Index. Science and Public Policy, 11(6):395–396, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:IAE

[Ano84e] Anonymous. Introduction to the ASEAN-EEC seminar on science and technology indicators and science policy. Science and Public Policy, 11(1):3, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:LFP

[Ano84f] Anonymous. Land, food and population. Science and Public Policy, 11(4):227–239, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:NAa

[Ano84g] Anonymous. Notes for authors. Science and Public Policy, 11(1): 69, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:NAb

[Ano84h] Anonymous. Notes for authors. Science and Public Policy, 11(5): 332, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:Pa

[Ano84i] Anonymous. Publications. Science and Public Policy, 11(1):67– 68, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:Pb

[Ano84j] Anonymous. Publications. Science and Public Policy, 11(2):117– 120, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:SCF

[Ano84k] Anonymous. Science — culturefree or culture-bound? Science and Public Policy, 11(3): 122, June 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:SPF

[Ano84l] Anonymous. The Science Policy Foundation (1964–1984). Science and Public Policy, 11(3):161– 172, June 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:SER

[Ano84m] Anonymous. Sf education resources. Science and Public Policy, 11(5):327–328, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1984:SWP

[Ano84n] Anonymous. Socialist world publications. Science and Public Policy, 11(4):255–256, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1985:CUS

[Ano85a] Anonymous. Continuing UK science bibliography. Science and Public Policy, 12(4):223, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1985:CC

[Ano85b] Anonymous. Copyright correction. Science and Public Policy, 12(5):296, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1985:EELa

[Ano85c] Anonymous. Eastern European literature. Science and Public Policy, 12(2):102–104, April 1985.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1985:EELb

[Ano85d] Anonymous. Eastern European literature. Science and Public Policy, 12(3):158–160, June 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1985:Ma

[Ano85e] Anonymous. Meetings. Science and Public Policy, 12(2):104– 105, April 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1985:Mb

[Ano85f] Anonymous. Meetings. Science and Public Policy, 12(3): co3, June 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1985:Mc

[Ano85g] Anonymous. Meetings. Science and Public Policy, 12(4): CO3, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1985:Md

[Ano85h] Anonymous. Meetings. Science and Public Policy, 12(5):295– 296, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1985:Me

[Ano85i] Anonymous. Meetings. Science and Public Policy, 12(6): CO3, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1985:PBI

[Ano85j] Anonymous. Partly baked ideas and problems. Science and Public Policy, 12(1):2, February 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1985:UL

[Ano85k] Anonymous. UK literature. Science and Public Policy, 12(2):99– 101, April 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1985:V

[Ano85l] Anonymous. Volume 12, 1985.

Science and Public Policy, 12(6): 359–360, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1986:CUS

[Ano86a] Anonymous. Continuing UK science bibliography. Science and Public Policy, 13(6):363–364, December 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1986:EELb

[Ano86b] Anonymous. East European literature. Science and Public Policy, 13(4):251–253, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1986:EELc

[Ano86c] Anonymous. East European literature. Science and Public Policy, 13(5):309–310, October 1986.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1986:EELa

[Ano86d] Anonymous. Eastern European literature. Science and Public Policy, 13(1):63–64, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1986:I

[Ano86e] Anonymous. Index. Science and Public Policy, 13(6):365–366, December 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1986:Ma

[Ano86f] Anonymous. Meetings. Science and Public Policy, 13(1): CO3, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1986:Mb

[Ano86g] Anonymous. Meetings. Science and Public Policy, 13(2): CO3, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1986:Mc

[Ano86h] Anonymous. Meetings. Science and Public Policy, 13(3): CO3, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1986:Md

[Ano86i] Anonymous. Meetings. Science and Public Policy, 13(4): co3, August 1986. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1986:Me

[Ano86j] Anonymous. Meetings. Science and Public Policy, 13(5): co3, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1986:Mf

[Ano86k] Anonymous. Meetings. Science and Public Policy, 13(6):co3, December 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1986:UL

[Ano86l] Anonymous. UK literature. Science and Public Policy, 13(4): 250, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:D

[Ano87a] Anonymous. Diary. Science and Public Policy, 14(2):CO3, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:EELa

[Ano87b] Anonymous. East European literature. Science and Public Policy, 14(2):109–110, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:EELb

[Ano87c] Anonymous. East European literature. Science and Public Policy, 14(3):175–176, June 1987. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:EELc

[Ano87d] Anonymous. East European literature. Science and Public Policy, 14(5):307–308, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:EELd

[Ano87e] Anonymous. East European literature. Science and Public Policy, 14(6):369–370, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:FMD

[Ano87f] Anonymous. Foundation’s Medal to Dr Joseph Needham. Science and Public Policy, 14(1):2, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:I

[Ano87g] Anonymous. Index. Science and Public Policy, 14(6):371, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:Ma

[Ano87h] Anonymous. Meetings. Science and Public Policy, 14(1): CO3, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:Mb

[Ano87i] Anonymous. Meetings. Science and Public Policy, 14(3):

CO3, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:Mc

[Ano87j] Anonymous. Meetings. Science and Public Policy, 14(4): CO3, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:Md

[Ano87k] Anonymous. Meetings. Science and Public Policy, 14(5): CO3, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:Me

[Ano87l] Anonymous. Meetings. Science and Public Policy, 14(6): CO3, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:NAa

[Ano87m] Anonymous. Notes for authors.

Science and Public Policy, 14(1): 42, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:NAb

[Ano87n] Anonymous. Notes for authors.

Science and Public Policy, 14(6): 346, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1987:UL

[Ano87o] Anonymous. UK literature. Science and Public Policy, 14(6):361, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:Da

[Ano88a] Anonymous. Diary. Science and Public Policy, 15(4):CO3, August 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:Db

[Ano88b] Anonymous. Diary. Science and Public Policy, 15(5):CO3, October 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:Dc

[Ano88c] Anonymous. Diary. Science and Public Policy, 15(6):CO3, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:EELa

[Ano88d] Anonymous. East European literature. Science and Public Policy, 15(2):127–128, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:EELb

[Ano88e] Anonymous. East European literature. Science and Public Policy, 15(4):273–275, August 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:EELc

[Ano88f] Anonymous. East European literature. Science and Public Policy, 15(6):437–438, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:FST

[Ano88g] Anonymous. A framework for science and technology policy analysis. Science and Public Policy, 15(4):194, August 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:IS

[Ano88h] Anonymous. Index, SPP 1988.

Science and Public Policy, 15(6): 439–440, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:Ma

[Ano88i] Anonymous. Meetings. Science and Public Policy, 15(1): CO3, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:Mb

[Ano88j] Anonymous. Meetings. Science and Public Policy, 15(2): co3, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:Mc

[Ano88k] Anonymous. Meetings. Science and Public Policy, 15(3): CO3, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:STP

[Ano88l] Anonymous. The Science, Technology and Public Policy Program, John F. Kennedy School of Government, Harvard University. Science and Public Policy, 15(5): 357–358, October 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:TN

[Ano88m] Anonymous. Technical note. Science and Public Policy, 15(6):366, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:ULa

[Ano88n] Anonymous. UK literature. Science and Public Policy, 15(3):191– 192, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1988:ULb

[Ano88o] Anonymous. UK literature. Science and Public Policy, 15(6):435– 436, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1989:EELa

[Ano89a] Anonymous. East European literature. Science and Public Policy, 16(1):62–64, February 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1989:EELb

[Ano89b] Anonymous. East European literature. Science and Public Policy, 16(2):135–136, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1989:EELc

[Ano89c] Anonymous. East European literature. Science and Public Policy, 16(4):259–260, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1989:EELd

[Ano89d] Anonymous. East European literature. Science and Public Policy, 16(5):317–320, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1989:EELe

[Ano89e] Anonymous. East European literature. Science and Public Policy, 16(6):381–382, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1989:IS

[Ano89f] Anonymous. Index, SPP 1989.

Science and Public Policy, 16(6): 383–384, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1989:Ma

[Ano89g] Anonymous. Meetings. Science and Public Policy, 16(1): CO3, February 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1989:Mb

[Ano89h] Anonymous. Meetings. Science and Public Policy, 16(2): co3, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1989:Mc

[Ano89i] Anonymous. Meetings. Science and Public Policy, 16(3): co3, June 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1989:Md

[Ano89j] Anonymous. Meetings. Science and Public Policy, 16(4): co3, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1989:Me

[Ano89k] Anonymous. Meetings. Science and Public Policy, 16(5): CO3, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1989:Mf

[Ano89l] Anonymous. Meetings. Science and Public Policy, 16(6):co3, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arnon:1980:FF

[Arn80] I. Arnon. From fellah to farmer. Science and Public Policy, 7(6): 443–447, December 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arnold:1983:LC

[Arn83] Arnold Arnold. Linking conflicts. Science and Public Policy, 10(6): 313–314, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arnold:1984:M

[Arn84] Arnold Arnold. Macroproblem.

Science and Public Policy, 11(1): 62–63, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arnold:1985:AIP

[Arn85] Arnold Arnold. Artificial intelligence — pro and con. Science and Public Policy, 12(2):97– 98, April 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arnold:1988:PWS

[Arn88a] Arnold Arnold. Playing at world suicide. Science and Public Policy, 15(1):71, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arnold:1988:PP

[Arn88b] Arnold Arnold. Proving a point.

Science and Public Policy, 15(3): 180–181, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arnold:1989:TMM

[Arn89] Arnold Arnold. Triumph of mind over matter. Science and Public Policy, 16(1):59–61, February 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ashby:1980:PIP

[Ash80] Lord Ashby. Pen-and-ink portrait of bernal. Science and Public Policy, 7(4):299–301, August 1980.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ashby:1982:DKR

[Ash82] Kenneth R. Ashby. Dr Kenneth R. Ashby analyses the problem of acid precipitation in Scandinavia, and presents a corrective to the CEGB view. Science and Public Policy, 9(2):106–110, April 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ashby:1983:CTE

[Ash83] Kenneth R. Ashby. Controlling the technocrat: Europe and America. Science and Public Policy, 10(4):202–204, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Baark:1988:HTI

[Baa88] Erik Baark. High technology innovation at the Chinese Academy of Sciences. Science and Public Policy, 15(2):81–90, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bahri:1985:EHL

[Bah85] A. Bahri. Expanded horizons: learning from another country. Science and Public Policy, 12(6): 344–347, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bailey:1984:JTT

[Bai84] Chris Bailey. Journey through time. Science and Public Policy, 11(5):330–331, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bajay:1982:ESI

[Baj82] S. V. Bajay. Economic and social impacts from a power system expansion program. Science and Public Policy, 9(4):179– 186, August 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Baker:1983:BCB

[Bak83] Rt Revd John Austin Baker. Beyond ‘the Church and the bomb’. Science and Public Policy, 10(4): 162–166, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Baker:1986:SDS

[Bak86] Dilys Baker. Sexual discrimination in science. Science and Public Policy, 13(5):302–303, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Baker:1987:AAS

[Bak87a] Dilys Baker. Are all aliens superintelligent? Science and Public Policy, 14(1):52–53, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Baker:1987:FV

[Bak87b] Dilys Baker. Feminist vehemence.

Science and Public Policy, 14(4): 239–240, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Balderrama:1988:ARP

[Bal88] Rafael Balderrama. Agricultural research priorities and farm modernization in Venezuela. Science and Public Policy, 15(3):163– 169, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barnaby:1983:LGC

[Bar83a] Frank Barnaby. Links in the glowing chain. Science and Public Policy, 10(2):92, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barnaby:1983:NWF

[Bar83b] Frank Barnaby. Nuclear war in focus. Science and Public Policy, 10 (6):311–313, December 1983. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barnaby:1983:UV

[Bar83c] Frank Barnaby. Unlimited values. Science and Public Policy, 10(1): 43–44, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barlow:1984:BDH

[Bar84] Kenneth Barlow. Bread, diet and health: a plea for policy. Science and Public Policy, 11(1):47– 51, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barrett:1987:ONN

[Bar87a] Dr Mark Barrett. Options for a non nuclear energy policy in the UK. Science and Public Policy, 14(5):292–296, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bartels:1987:HER

[Bar87b] Dr Ditta Bartels. Human embryos as research material. Science and Public Policy, 14(3):139– 144, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bauer:1987:MSS

[Bau87] Henry H. Bauer. Mediating science and society. Science and Public Policy, 14(2):113–114, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bazelon:1984:ILT

[Baz84] David L. Bazelon. Interaction of law and technology. Science and Public Policy, 11(4):240– 242, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bundock:1985:FWF

[BBGM85] J. B. Bundock, D. Burk, J. R.

Graham, and P. J. Morin. Fluorides, water fluoridation, cancer and genetic diseases. Science and Public Policy, 12(1):36– 46, February 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1980:IR

[BC80a] M. R. Brett-Crowther. The imbroglio of risk. Science and Public Policy, 7(5):306–307, October 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1980:IBS

[BC80b] M. R. Brett-Crowther. Interactions between science and government in China since 1949. Science and Public Policy, 7(2):98– 119, April 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1980:UDM

[BC80c] M. R. Brett-Crowther. Uncertain decision making on environmental problems. Science and Public Policy, 7(5):377–393, October 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

B-C:1981:FRE

[BC81a] M. R. B-C. Fells rises in the energy debate. Science and Public Policy, 8(1):2–4, February 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1981:LZQ

[BC81b] M. R. Brett-Crowther. Lord Zuckerman’s question. Science and Public Policy, 8(5):342–344, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1981:MMP

[BC81c] M. R. Brett-Crowther. Moral and material policies. Science and Public Policy, 8(2):157–162, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1981:SP

[BC81d] M. R. Brett-Crowther. Severed pieces. Science and Public Policy, 8(3):170–173, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1981:SBC

[BC81e] M. R. Brett-Crowther. The spruce budworm controversy in Nova Scotia. Science and Public Policy, 8(1):55–76, February 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1981:CI

[BC81f] Michael Brett-Crowther. Circular information. Science and Public Policy, 8(4):258–261, August 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1981:EE

[BC81g] Michael Brett-Crowther. Extrapolating and escaping. Science and Public Policy, 8(2):82– 84, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1981:VS

[BC81h] Michael Brett-Crowther. Values in science. Science and Public Policy, 8(6):422–425, December 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1982:LAN

[BC82a] M. R. Brett-Crowther. Life after nuclear war: the economic and social impacts of nuclear attacks on the United States; the nuclear power decisions: British policies, 1953–78; nuclear power and legal advocacy: the environmentalists and the courts; energy in a finite world: paths to a sustainable future; oil and security in the Arabian Gulf; the economic development of the United Arab Emirates; building a sustainable society; the God that limps. Science and Public Policy, 9(4):221– 223, August 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1982:PSP

[BC82b] M. R. Brett-Crowther. The philosophy of science and its possible effects. Science and Public Policy, 9(1):35–47, February 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1983:S

[BC83a] Lucy A. Brett-Crowther. Statecraft. Science and Public Policy, 10(5):261–262, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1983:CPR

[BC83b] M. R. Brett-Crowther. Chinese puzzles reconstructed. Science and Public Policy, 10(2):93– 96, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1983:C

[BC83c] M. R. Brett-Crowther. Compendium. Science and Public Policy, 10(3):157–158, June 1983.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1983:II

[BC83d] M. R. Brett-Crowther. Indian images. Science and Public Policy, 10(1):40–42, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1983:OOI

[BC83e] M. R. Brett-Crowther. Occidenting orientals without incident. Science and Public Policy, 10(5):260– 261, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1983:TWC

[BC83f] M. R. Brett-Crowther. Technology, war and conscience. Science and Public Policy, 10(4):167– 176, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1983:VMB

[BC83g] M. R. Brett-Crowther. The value of a moral Baedeker: M. R. Brett-Crowther. Science and Public Policy, 10(4):212–215, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1983:WW

[BC83h] M. R. Brett-Crowther. War on want. Science and Public Policy, 10(5):259–260, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1984:DE

[BC84] M. R. Brett-Crowther. Dynamics of the establishment. Science and Public Policy, 11(2):112– 116, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bell:1982:BSS

[Bel82] E. A. Bell. Botany: the science of survival. Science and Public Policy, 9(4):170–178, August 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bellany:1986:DAC

[Bel86] Ian Bellany. Is disinterested arms control attainable? Science and Public Policy, 13(2): 120, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Benfey:1987:SIM

[Ben87] Otto Theodor Benfey. Scientific instruments, Meiji scholars and the relevance of the Orient for modern science. Science and Public Policy, 14(6):329–335, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bertrand:1980:PRS

[Ber80] Anson R. Bertrand. Public responsibilities of the scientist. Science and Public Policy, 7(3):216– 218, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bernal:1989:ATF

[Ber89] J. D. Bernal. After twenty-five years. Science and Public Policy, 16(3):143–151, June 1989. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Breakwell:1985:SSA

[BFSL85] Dr Glynis M. Breakwell, Dr Chris Fife-Schaw, and Professor Terence Lee. Survey of student attitudes to technology. Science and Public Policy, 12(6):337–340, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bundock:1982:FWF

[BGM82] J. B. Bundock, J. R. Graham, and P. J. Morin. Fluorides, water fluoridation and environmental quality. Science and Public Policy, 9 (3):130–142, June 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bhatt:1982:FIT

[Bha82] V. V. Bhatt. Financial institutions and technology policy. Science and Public Policy, 9(4):203– 209, August 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Biswas:1980:CSP

[Bis80] Sanat Biswas. Case study on the popularization of science and technology. Science and Public Policy, 7(2):135–141, April 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Blomley:1983:PNA

[Blo83] Peter Blomley. Prospects for nuclear arms reductions. Science and Public Policy, 10(2):50– 55, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Blume:1985:SDA

[Blu85] Stuart Blume. Science dynamics at Amsterdam. Science and Public Policy, 12(6):351–352, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barbier:1985:DWC

[BM85a] Ed Barbier and Nick Mahoney. Do workers’ co-operatives innovate in the Third World? Science and Public Policy, 12(2):67– 71, April 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bowling:1985:SMD

[BM85b] Jill Bowling and Brian Martin. Science: a masculine disorder? Science and Public Policy, 12(6): 308–316, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bowonder:1988:MII

[BM88a] B. Bowonder and T. Miyake. Measuring innovativeness of an industry: an analysis of the electronics industry in India, Japan and Korea. Science and Public Policy, 15(5):279–303, October 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bowonder:1988:MTI

[BM88b] Dr B. Bowonder and Dr T.

Miyake. Measurement of technology at industry level: a case study of the steel industry in India and Japan. Science and Public Policy, 15(4):249–269, August 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bondi:1980:SPS

[Bon80] Hermann Bondi. Scientists in public service. Science and Public Policy, 7(4):234–241, August 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bondi:1981:SWY

[Bon81] Sir Hermann Bondi. Scintillating war years. Science and Public Policy, 8(2):164–165, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bondi:1984:PRD

[Bon84] Sir Herman Bondi. Policies of R and D. Science and Public Policy, 11(4):184–187, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bondi:1986:GBL

[Bon86a] Herman Bondi. Great Britain Ltd (R&D Division). Science and Public Policy, 13(3):173–174, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bondi:1986:TE

[Bon86b] Hermann Bondi. Technology and employment. Science and Public Policy, 13(5):256–258, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Borlin:1986:UNR

[Bör86] Max Börlin. Use of natural resources and pollution of the environment. Science and Public Policy, 13(4):230–235, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bowonder:1980:EMT

[Bow80] B. Bowonder. Environmental management and the Third World. Science and Public Policy, 7(3):185–198, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bowles:1984:RDM

[Bow84a] J. R. Bowles. Research and development: measures of input. Science and Public Policy, 11(1):25– 28, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bowonder:1984:CTE

[Bow84b] B. Bowonder. Catastrophe theory and environmental changes. Science and Public Policy, 11(2): 94–99, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boyle:1983:WOA

[Boy83] John E. Whiteford Boyle. The welcome ordeal: an aspect of Iran’s westernization drive 1957– 62. Science and Public Policy, 10 (5):240–243, October 1983. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boyle:1985:TE

[Boy85] Charles Boyle. Text on the emerging. Science and Public Policy, 12(4):226–227, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boyle:1986:CSF

[Boy86a] Charles Boyle. Classification of scientific fields. Science and Public Policy, 13(1):58–59, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boyle:1986:MPD

[Boy86b] Charles Boyle. My public doesn’t understand me. Science and Public Policy, 13(3):175–176, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boyle:1987:NO

[Boy87] Charles Boyle. New for old. Science and Public Policy, 14(6):353– 354, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boyle:1988:SB

[Boy88] Charles Boyle. Surprising bland.

Science and Public Policy, 15(5): 361–362, October 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bowonder:1986:OEI

[BR86] B. Bowonder and K. V. Ramana. Occupational environment in India: policy implications. Science and Public Policy, 13(6):321–332, December 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brunat:1989:LUI

[BR89] Eric Brunat and Bernard Reverdy. Linking university and industrial research in France. Science and Public Policy, 16(5):283–293, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bradbury:1980:DAV

[Bra80] Farel Bradbury. Democratic accountability and the votes for nuclear energy. Science and Public Policy, 7(5):346–355, October 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bradbury:1981:AVA

[Bra81a] Farel Bradbury. The abolition of value added tax. Science and Public Policy, 8(3):205– 207, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bradbury:1981:IIM

[Bra81b] R. H. Bradbury. The international impact of microelectronics. Science and Public Policy, 8(4):281– 290, August 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brazier:1984:AG

[Bra84] Paul Brazier. Anti-gravity. Science and Public Policy, 11(5):275– 278, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Britten:1982:SBE

[Bri82] Stewart Britten. Stewart Britten emphasises psychological factors in nuclear weapons deployment. Science and Public Policy, 9(2):71–80, April 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brockman:1983:ALC

[Bro83a] John R. Brockman. Academic library costs. Science and Public Policy, 10(3):153–154, June 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brookman:1983:DSP

[Bro83b] F. Henry Brookman. The development of science policy in The Netherlands, 1945–1975. Science and Public Policy, 10(3):134– 141, June 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brookman:1983:MKH

[Bro83c] F. Henry Brookman. Models of know-how and technology transfer in the international order. Science and Public Policy, 10(5):229– 239, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brown:1983:CH

[Bro83d] Colin Brown. Crash hot. Science and Public Policy, 10(2):96– 97, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brown:1983:E

[Bro83e] Colin Brown. Into Europe. Science and Public Policy, 10(1):45– 46, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brown:1983:PTD

[Bro83f] Colin Brown. Politics and technology of defence. Science and Public Policy, 10(1):46–47, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brown:1983:WVC

[Bro83g] Colin D. Brown. Western values: a case for re-appraisal. Science and Public Policy, 10(5): 255, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brown:1984:SBI

[Bro84] Colin D. Brown. Svedenstierna and the British iron industry, 1802-03. Science and Public Policy, 11(2):106–108, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brooks:1986:RLE

[Bro86] Chris W. Brooks. Role of local employment initiatives. Science and Public Policy, 13(4):204– 206, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brunner:1985:CCD

[Bru85a] John Brunner. Childless couples and delinquent children. Science and Public Policy, 12(3):149– 152, June 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruyn:1985:TWS

[Bru85b] Severyn T. Bruyn. The transition to worker self management in the United States. Science and Public Policy, 12(2):79–83, April 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1982:DL

[BS82a] David Bridgeman-Sutton. Dirty but legal? Science and Public Policy, 9(5):273–274, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1982:SDG

[BS82b] David Bridgeman-Sutton. Scientists and decisionmaking: Guidedogs or lapdogs? Science and Public Policy, 9(6):330–332, December 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1983:TNL

[BS83a] D. Bridgeman-Sutton. Technology: need for the long view. Science and Public Policy, 10(6):310– 311, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1983:HIC

[BS83b] David Bridgeman-Sutton. Hot issues at the core. Science and Public Policy, 10(6):314–315, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1983:HVM

[BS83c] David Bridgeman-Sutton. How very much things will alter. Science and Public Policy, 10(2):65– 70, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1983:LTM

[BS83d] David Bridgeman-Sutton. Location, technology and management. Science and Public Policy, 10(3): 154–157, June 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1983:MUS

[BS83e] David Bridgeman-Sutton. Malaysia: a unique study in consumerism. Science and Public Policy, 10(2): 86–88, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1983:NAR

[BS83f] David Bridgeman-Sutton. North American risk management. Science and Public Policy, 10(4):205– 206, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1983:OU

[BS83g] David Bridgeman-Sutton. Onwards and upwards. Science and Public Policy, 10(5):254–255, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1983:SSR

[BS83h] David Bridgeman-Sutton. Setting sun, rising sun: British and Japanese management. Science and Public Policy, 10(4):207– 208, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1983:TT

[BS83i] David Bridgeman-Sutton. Titans and technology. Science and Public Policy, 10(5):253–254, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgemann-Sutton:1983:PRP

[BS83j] David Bridgemann-Sutton. Population, resources and policy. Science and Public Policy, 10(1):47– 48, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1984:MM

[BS84a] D. Bridgeman-Sutton. Miracle without myth. Science and Public Policy, 11(2):110–111, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1984:BE

[BS84b] David Bridgeman-Sutton. Brazilian enterprise. Science and Public Policy, 11(1):63–64, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bridgeman-Sutton:1984:RCD

[BS84c] David Bridgeman-Sutton. Reading the crowded display. Science and Public Policy, 11(2):111– 112, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Budworth:1989:IAE

[Bud89a] D. W. Budworth. Intangible assets evaluated. Science and Public Policy, 16(6):372–373, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Budworth:1989:TRE

[Bud89b] Dr D. W. Budworth. Technology recognisable in economics. Science and Public Policy, 16(1):56– 57, February 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bueche:1980:IUF

[Bue80] Arthur M. Bueche. Innovation in the USA: 2. fears of a longterm decline are exaggerated. Science and Public Policy, 7(3):166– 170, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bullough:1981:NCC

[Bul81] K. Bullough. The new colony:

a contribution to peace. Science and Public Policy, 8(5):361– 375, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bullough:1983:MPR

[Bul83] K. Bullough. Military perception and reality: Thoughts on Lord Carver’s Policy for Peace. Science and Public Policy, 10(1):34– 37, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Burkitt:1981:DP

[Bur81] D. P. Burkitt. Disease prevention.

Science and Public Policy, 8(6): 433–443, December 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brookman:1981:DIN

[BV81] F. H. Brookman and A. J. C. Vierling. Dutch involvement in nuclear armament. Science and Public Policy, 8(5):389–395, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cahill:1983:WNH

[Cah83] Eileen J. Cahill. Woman’s nature, human nature. Science and Public Policy, 10(1):39–40, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cameron:1982:ECA

[Cam82] N. Cameron. Estimation of chronological age in children. Science and Public Policy, 9(1):20– 27, February 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Campbell:1983:CPD

[Cam83] Sir Colin Campbell. Commodity prices and development. Science and Public Policy, 10(3):152– 153, June 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Campbell:1984:FTH

[Cam84] Sir Colin Campbell. Further thoughts on helping the Third World. Science and Public Policy, 11(1):52–53, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Campbell:1986:JTC

[Cam86] Dr Nigel Campbell. Japan thurowly considered. Science and Public Policy, 13(3):177–178, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Carley:1986:PAT

[Car86] Michael J. Carley. A policy approach to technology assessment: values and the future. Science and Public Policy, 13(2):77– 82, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cernosek:1980:EO

[Cer80] Marilyn G. Cernosek. Economic outlook for the 1980s. Science and Public Policy, 7(1):2– 9, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Champion:1988:MP

[Cha88] Lynn Champion. Master of paradox. Science and Public Policy, 15(1):70–71, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Chanaron:1989:FSP

[Cha89] J. J. Chanaron. French science policy and local high tech industries. Science and Public Policy, 16(1):19–26, February 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Chen:1988:CPL

[Che88] Ching-Shing Chen. Communist policy legislation on S&T — the model of the People’s Republic of China. Science and Public Policy, 15(3):131–136, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Chubin:1987:DRP

[Chu87] Daryl E. Chubin. Designing research program evaluations: a science studies approach. Science and Public Policy, 14(2):82– 90, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Clark:1988:NAL

[Cla88] Norman Clark. New academic languages required. Science and Public Policy, 15(3):183–184, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cliff:1980:NRC

[CO80] K. D. Cliff and M. C. O’Riordan. Natural radioactivity in the countries of the European Community. Science and Public Policy, 7(4): 281–288, August 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cohen:1981:PCC

[Coh81] E. A. Cohen. The postconcentration camp syndrome: a disaster syndrome. Science and Public Policy, 8(3):239–246, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Coleman:1981:CLE

[Col81a] Alice Coleman. City limits: Emerging constraints on urban growth. Science and Public Policy, 8(3):250, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Coleman:1981:LUL

[Col81b] Alice Coleman. Land use and leisure demands. Science and Public Policy, 8(1):32–39, February 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Coleman:1981:OO

[Col81c] Alice Coleman. Opportunity orientation. Science and Public Policy, 8(3):208–216, June 1981. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Coleman:1982:WTD

[Col82a] Alice Coleman. Woods and trees distinguished. Science and Public Policy, 9(5):276–277, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Coleman:1982:S

[Col82b] Alice M. Coleman. Spaces. Science and Public Policy, 9(4):220– 221, August 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Coleman:1983:BRB

[Col83] Alice M. Coleman. Book review:

Problems and Planning in Third World cities, ed Michael Pacione, Croom Helm, London, 1981, 304 pages, £12.95. Science and Public Policy, 10(4):215, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Collingridge:1984:TNA

[Col84] David Collingridge. Technology in the nuclear arms race. Science and Public Policy, 11(3):173– 177, June 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Colombo:1986:ETP

[Col86] Umberto Colombo. European technology perspective. Science and Public Policy, 13(4):239–242, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Coleman:1988:HR

[Col88a] Alice Coleman. High rise. Science and Public Policy, 15(2):99– 107, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Collins:1988:MST

[Col88b] Eileen L. Collins. Meeting the scientific and technical staffing requirements of the American economy. Science and Public Policy, 15(5):335–342, October 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cooke:1983:RMN

[Coo83] Rev C. K. Hamel Cooke. Religion and medicine — a new partnership. Science and Public Policy, 10 (3):142–145, June 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cooney:1984:PTT

[Coo84] Sean Cooney. Progress through technology: need for a new assessment. Science and Public Policy, 11(1):29–36, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cottrell:1986:GRD

[Cot86] Alan Cottrell. Government research and development reorganisation in the UK in 1972. Science and Public Policy, 13(1):55– 57, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cox:1983:UAR

[Cox83] Sir Gordon Cox. UK agricultural research at the crossroads. Science and Public Policy, 10(4):184– 191, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cozzens:1987:ERE

[Coz87] Susan E. Cozzens. Expert review in evaluating programs. Science and Public Policy, 14(2):71– 81, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Critchley:1981:TPS

[Cri81]  O. H. Critchley.  Technolog-

ical progress, safety, and the guardian role of inspection. Science and Public Policy, 8(4):291– 307, August 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cropley:1984:LEA

[Cro84] Arthur Cropley. Lifelong education in an age of electronic revolution. Science and Public Policy, 11 (6):369–373, December 1984. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cromwell:1989:GPA

[Cro89a] Godfrey Cromwell. Government policy and alternative strategies for appropriate technology choice. Science and Public Policy, 16(4): 202–210, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Crow:1989:TDJ

[Cro89b] Dr Michael Crow. Technology development in Japan and the United States: Lessons from the high-temperature superconductivity race. Science and Public Policy, 16(6):322–344, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Csizmadia:1982:DIS

[Csi82] I. G. Csizmadia. The dialectical interaction of science, technology, economy and politics: a quest for rational research policies. Science and Public Policy, 9(5):255– 261, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cummings:1987:LSB

[Cum87] William K. Cummings. Life science and biotechnology R&D in Japan. Science and Public Policy, 14(6):317–328, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Curie:1981:DON

[Cur81] Adam Curie. Do our nuclear weapons protect us more than they imperil us? Science and Public Policy, 8(5):350–360, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Davies:1982:DPA

[Dav82] Jean Davies. Dangers in the policy against home confinements. Science and Public Policy, 9(6):286– 291, December 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Davis:1984:BRB

[Dav84] Hazel Davis. Book review: How to Organise Information: a Manager’s Guide to Techniques and Sources, with a Checklist for Secretaries and Assistants by K. G. B. Bakewell Gower, Aldershot, UK, 1984, 225pp. Science and Public Policy, 11(6):393, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Davidson:1987:LPU

[Dav87a] Dr Frank P. Davidson. Large projects and the US presidential race. Science and Public Policy, 14 (6):365–368, December 1987. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Davies:1987:CC

[Dav87b] T. D. Davies. Climatic change.

Science and Public Policy, 14(3): 171–174, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Davison:1987:PMD

[Dav87c] D. J. Davison. Project management diagnosis. Science and Public Policy, 14(5):300–301, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Davis:1988:STP

[Dav88] Charles H. Davis. Science, technology and policy in Quebec. Science and Public Policy, 15(1):26– 34, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dunmore:1982:FEC

[DB82] Jane Dunmore and Ron Barnes. First European Conference on Acid Rain Göteborg, Sweden, May 1981. Science and Public Policy, 9(2):104–106, April 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Demetris:1985:DCR

[Dem85] Dr Deniozos Demetris. Developing country research: experience from Greece. Science and Public Policy, 12(6):341–343, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Derrick:1985:TCC

[Der85] Paul Derrick. Towards a consensus on common ownership. Science and Public Policy, 12(2):55– 60, April 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

DeWitt:1982:DHD

[DeW82] Hugh E. DeWitt. Dr Hugh E.

Dewitt contemplates the nuclear arms race from within an American weapons laboratory. Science and Public Policy, 9(2):58– 63, April 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dey:1984:WRF

[Dey84] Dr Jennie Dey. Women in rice farming systems. Science and Public Policy, 11(4):201–218, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dufour:1988:DCS

[DG88] Paul Dufour and Yves Gingras.

Development of Canadian science and technology policy. Science and Public Policy, 15(1):13–18, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dhar:1984:CCI

[Dha84] Dr M. M. Dhar. A comparison of cancer incidence. Science and Public Policy, 11(3): 180, June 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dickson:1986:OSO

[Dic86] Dr Keith Dickson. Old stuff but original theme. Science and Public Policy, 13(3):181–182, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1988:DB

[Div88] Colin Divall. Difficult birth. Science and Public Policy, 15(5):362– 364, October 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1989:IS

[Div89a] Colin Divall. Ideological struggles. Science and Public Policy, 16(6): 376–378, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Divall:1989:PPS

[Div89b] Colin Divall. Physicist, philosopher and statesman. Science and Public Policy, 16(5):310–311, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

delaMothe:1988:ASE

[dlM88a] John de la Mothe. Always someone else’s fault. Science and Public Policy, 15(1):66–67, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

delaMothe:1988:CST

[dlM88b] John de la Mothe. Canadian science and technology policy. Science and Public Policy, 15(1):2, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

delaMothe:1988:CIE

[dlM88c] John R. de la Mothe. Changing industrial economy. Science and Public Policy, 15(2):120– 121, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

delaMothe:1988:AE

[dlM88d] John R. de la Mothe. From apathy to engagement. Science and Public Policy, 15(6):432–433, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

delaMothe:1988:TPF

[dlM88e] John R. de la Mothe. Technology policy and the Freeman kindled fire. Science and Public Policy, 15 (2):119–120, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Guimaraes:1989:TPN

[dMSG89] Fabio Celso de Macedo Soares Guimarães.

Technology policy in newly industrialised countries: a Brazilian perspective. Science and Public Policy, 16(3):167–175, June 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dodd:1989:RB

[Dod89a] W. A. Dodd. Readable and balanced. Science and Public Policy, 16(4):254–255, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dodgson:1989:RTP

[Dod89b] Dr Mark Dodgson. Research and technology policy in Australia: legitimacy in intervention. Science and Public Policy, 16(3):159– 166, June 1989. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Donnet:1980:RS

[Don80] J. B. Donnet. La responsibilité de scientific. Science and Public Policy, 7(2):141–145, April 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Donohew:1984:SIA

[Don84] Lewis Donohew. Some implications of automaticity and arousal for the mass media. Science and Public Policy, 11(6):388–390, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dore:1986:WWJ

[Dor86] Ron Dore. Where will the Japanese Nobel Prizes come from? Science and Public Policy, 13(6): 347–361, December 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

deReuck:1987:TU

[dR87] Anthony de Reuck. Totally unconvincing. Science and Public Policy, 14(6):355–357, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

deReuck:1988:ECV

[dR88] Anthony de Reuck. Extraordinary clarity of vision. Science and Public Policy, 15(2):122– 123, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dror:1980:WEL

[Dro80] Yehezkel Dror. The work environment of leaders. Science and Public Policy, 7(3):220–222, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

deSilva:1989:SPS

[dS89] M. A. T. de Silva. Science planning: The Sri Lankan perspective. Science and Public Policy, 16(6): 367–371, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Price:1989:SS

[dSP89] D. J. de Solla Price. The science of science. Science and Public Policy, 16(3):152–158, June 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dobrov:1987:SRU

[DST87] Gennady M. Dobrov, Anatoly A.

Saveliev, and Vladimir E. Tonkal. Studying research utilization in the Ukraine. Science and Public Policy, 14(4):231–235, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Duckworth:1983:DBT

[Duc83] W. E. Duckworth. Doing better together: the personal factor in innovation. Science and Public Policy, 10(2):77–79, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ducharme:1988:UFE

[Duc88] Louis Marc Ducharme. From unemployment to full employment. Science and Public Policy, 15(1): 67–68, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Duffy:1984:E

[Duf84] Michael Duffy. Education 2000? Science and Public Policy, 11(6): 364–368, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Edgerton:1989:LOC

[Edg89] David Edgerton. Long overdue correction of misapprehensions. Science and Public Policy, 16(2):131–132, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Eisemon:1985:TSA

[EDR85] Thomas O. Eisemon, Charles H. Davis, and Eva-Marie Rathgeber. Transplantation of science to Anglophone and Francophone Africa. Science and Public Policy, 12(4): 191–202, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ega:1988:EWB

[Ega88] L. Alegwu Ega. Experience of World Bank assistance in agriculture in Ayangba, Nigeria. Science and Public Policy, 15(2):91– 98, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Egerton:1989:SMD

[Ege89] Dr David Egerton. Staggering missions and disarming innocence. Science and Public Policy, 16(5): 313–314, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ehrlich:1980:DSS

[Ehr80] Paul R. Ehrlich. Diversity and the steady state. Science and Public Policy, 7(4):252–264, August 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Einsiedel:1984:MAN

[Ein84] Edna F. Einsiedel. Media agendas and the new technologies: Nurturing the information gap. Science and Public Policy, 11(6):378–380, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Eisemon:1984:EED

[Eis84] Thomas Eisemon. Educational expansion and the development of science in Kenya. Science and Public Policy, 11(2):70–76, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ekins:1986:HW

[Eki86a] Paul Ekins. Health as wealth. Science and Public Policy, 13(4):194– 195, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ekins:1986:IEP

[Eki86b] Paul Ekins. Indicators of economic progress. Science and Public Policy, 13(4):216–217, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Elliott:1987:RET

[Ell87] Dr David Elliott. Renewable energy and the tidal power options. Science and Public Policy, 14(5): 257–262, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Emmelin:1980:SBW

[Emm80] Lars Emmelin. The significance of the Belgrade workshop, the Helsinki regional meeting, and the intergovernmental conference, Tbilisi. Science and Public Policy, 7(1):30–35, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ekbote:1980:FSK

[EMR^{+}80] P. D. Ekbote, S. Mohan, P. K. Rohatgi, H. Sreemoolanathan, S. Suresh Kumar, and R. N. Yadava. A forecasting study of Kerala. Science and Public Policy, 7 (3):198–207, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Eng:1980:RR

[Eng80] Sir Frederick Warner F. Eng. Risk and response. Science and Public Policy, 7(6):402–406, December 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fabian:1984:OII

[Fab84] Dr Yvan Fabian. The OECD international S and T indicators system. Science and Public Policy, 11 (1):4–6, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fairbairn:1982:IIT

[Fai82] David R. Fairbairn. The impact of information technology on employment. Science and Public Policy, 9(5):236–239, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Falk:1984:GST

[Fal84] Charles E. Falk. Guidelines for science and technology indicators projects. Science and Public Policy, 11(1):37–39, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Famoriyo:1986:ARA

[Fam86] O. A. Famoriyo. Agricultural research administration in Nigeria. Science and Public Policy, 13(5): 295–300, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fasick:1984:EPC

[Fas84] Adele Fasick. Education policy and the communication gap. Science and Public Policy, 11(6):384– 385, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fells:1981:NEW

[Fel81] Ian Fells. Nuclear energy — the way ahead. Science and Public Policy, 8(3):217–226, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fels:1988:EMF

[Fel88] Xavier Fels. Eureka, a model for the future. Science and Public Policy, 15(6):373–375, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ferguson:1981:SI

[Fer81] R. A. D. Ferguson. Striking inefficiencies. Science and Public Policy, 8(5):405–407, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ferguson:1982:REE

[Fer82] R. A. D. Ferguson. Risk estimation and evaluation. Science and Public Policy, 9(5):251– 254, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Few:1980:IWC

[Few80] Arthur A. Few, Jr. The implications of a widespread conversion to biomass-based fuels. Science and Public Policy, 7(4):264– 275, August 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ford:1986:COS

[FGSA86] Glyn Ford, Ken Green, Fred Steward, and Salma Ahmad. Coordinating strategic research in Britain. Science and Public Policy, 13(5):269–274, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fitzpatrick:1985:CPP

[Fit85a] Mary Anne Fitzpatrick. Communications past, present and future. Science and Public Policy, 12(1): 48, February 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fitzpatrick:1985:HTC

[Fit85b] Mary Anne Fitzpatrick. High technology culture. Science and Public Policy, 12(1):47, February 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fitzpatrick:1985:SIW

[Fit85c] Mary Anne Fitzpatrick. Say it or write it. Science and Public Policy, 12(1):48–49, February 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forje:1985:BPE

[For85a] John W. Forje. Booknotes: politics, economy, ecology and technology in the Third World. Science and Public Policy, 12(4):224– 225, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forje:1985:WWA

[For85b] John W. Forje. Which way Africa?: The shameful waiting for exogenous initiatives. Science and Public Policy, 12(4):162– 163, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forje:1985:AIU

[For85c] Lema Catherine Forje. Africa’s inability to use transnational firms for development. Science and Public Policy, 12(4):203–206, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forje:1986:CTP

[For86a] John W. Forje. Coming to terms with population explosion. Science and Public Policy, 13(5):304– 305, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forje:1986:ES

[For86b] John W. Forje. European sandwich. Science and Public Policy, 13(3):178, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forje:1986:TDS

[For86c] John W. Forje. Two decades of science and technology in Africa. Science and Public Policy, 13(2): 89–96, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forje:1986:WGB

[For86d] John W. Forje. Widening gap between poverty and affluence. Science and Public Policy, 13(5):306– 307, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ford:1987:SSN

[For87] Glyn Ford. Science and scientists neglected. Science and Public Policy, 14(2):115–116, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ford:1988:CAS

[For88a] Glyn Ford. Campaign against state terrorism. Science and Public Policy, 15(3):187–188, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ford:1988:ETC

[For88b] Glyn Ford. European technological choices. Science and Public Policy, 15(6):367–372, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ford:1988:FP

[For88c] Glyn Ford. Future promise. Science and Public Policy, 15(1):68– 69, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ford:1988:NFC

[For88d] Glyn Ford. New found confidence.

Science and Public Policy, 15(6): 431–432, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forester:1988:NM

[For88e] Tom Forester. New materials. Science and Public Policy, 15(4):271– 272, August 1988. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Ford:1989:DEJ

[For89] Chris Ford. Demand for equality and justice. Science and Public Policy, 16(3):198–199, June 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Freedman:1981:RPF

[Fre81] Ron Freedman. Reentry programs for female scientists. Science and Public Policy, 8(3):249, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Freedman:1984:PSP

[Fre84] Ron Freedman. A politician on science policy. Science and Public Policy, 11(2):109–110, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Freeman:1988:SA

[Fre88] Hugh Freeman. Sound advice. Science and Public Policy, 15(2):123– 124, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Frookman:1984:CP

[Fro84] F. Henry Frookman. Centre and periphery. Science and Public Policy, 11(1):57–58, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fryers:1980:HGN

[Fry80a] Gordon Fryers. The Hoyle guide to nuclear power. Science and Public Policy, 7(5):398, October 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fryers:1980:II

[Fry80b] Gordon Fryers. Innovation and investment. Science and Public Policy, 7(1):67–69, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fryers:1982:DP

[Fry82] Gordon Fryers. Defining the problem. Science and Public Policy, 9 (5):276, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forestier-Walker:1982:ITP

[FW82] C. O. Forestier-Walker. Integrated technology projects for rural communities. Science and Public Policy, 9(5):226–235, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

G:1984:SJA

[G.84] M. G. Sir John Adams CMG, FRS. Science and Public Policy, 11(3):123–124, June 1984. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

G:1985:BNS

[G.85a] M. G. Britain’s next science and technology policy? Science and Public Policy, 12(6):298, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

G:1985:LBM

[G.85b] M. G. Lord Brown of Machrinhanish PC, MBE. Science and Public Policy, 12(3):157, June 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

G:1985:SI

[G85c] M. G. Special issue. Science and Public Policy, 12(2): 54, April 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Galbraith:1987:MPT

[Gal87] Professor John Kenneth Galbraith. Military power: tension as a servant, arms control as an illusion. Science and Public Policy, 14(1):19–22, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Garfield:1983:MST

[Gar83] Eugene Garfield. Mapping science in the Third World. Science and Public Policy, 10(3):112– 127, June 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gardiner:1987:SBI

[Gar87] Paul Gardiner. Some British innovations and what happened to them. Science and Public Policy, 14(3):177, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Green:1980:AR

[GB80] C. H. Green and R. A. Brown. The acceptability of risk. Science and Public Policy, 7(5):307– 318, October 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goncalves:1987:STP

[GC87] Fernando Gonçalves and J. M. G.

Caraça. Science and technology in Portugal: a perspective. Science and Public Policy, 14(3):155– 158, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Greenland:1984:RA

[GD84] Colin Greenland and Roger Dean.

Radical architecture. Science and Public Policy, 11(5):281–283, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Geis:1986:ST

[Gei86] Richard E. Geis. In search of truth. Science and Public Policy, 13(4):248, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gentle:1984:EFU

[Gen84] Mary Gentle. Exploring the food Utopia. Science and Public Policy, 11(5):329, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Georghiou:1986:EFC

[Geo86] Dr Luke Georghiou. Europe — a fragmented continent? Science and Public Policy, 13(2):117– 118, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Georghiou:1989:FF

[Geo89] Luke Georghiou. Foresight not forecasting. Science and Public Policy, 16(6):379–380, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ghee:1982:STD

[Ghe82] Ang How Ghee. Science and technology for development in ASEAN countries. Science and Public Policy, 9(6):297–301, December 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Giarini:1980:EDE

[Gia80] Orio Giarini. Economic development, economic theory, and risks. Science and Public Policy, 7(5): 336–345, October 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Giarini:1986:CUD

[Gia86a] Orio Giarini. Challenge of uncertainty: a dialogue. Science and Public Policy, 13(4):236–238, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Giarini:1986:CAS

[Gia86b] Orio Giarini. Coming of age of the service economy. Science and Public Policy, 13(4):209– 215, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Giarini:1986:DPA

[Gia86c] Orio Giarini. Developing productive activities for the wealth of nations in the service economy. Science and Public Policy, 13(4):221– 229, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gilland:1983:D

[Gil83a] Bernard Gilland. Demythologizing. Science and Public Policy, 10 (5):257, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gilland:1983:EE

[Gil83b] Bernard Gilland. Energy and equity. Science and Public Policy, 10(4):210–211, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gilland:1983:WPF

[Gil83c] Bernard Gilland. World population and food supply: a realist view. Science and Public Policy, 10(3):149–151, June 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gilland:1984:CP

[Gil84] Bernard Gilland. CO 2 in perspective. Science and Public Policy, 11(2):90–93, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1989:AFY

[GM89a] Maurice Goldsmith and Alan Mackay. After fifty years. Science and Public Policy, 16(3):

138, June 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1989:CB

[GM89b] Maurice Goldsmith and Alan Mackay. In commemoration of bernal. Science and Public Policy, 16(3):139–142, June 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1980:WPS

[Gol80] Maurice Goldsmith. World policy and self-reliance. Science and Public Policy, 7(2):82, April 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1981:C

[Gol81a] Maurice Goldsmith. Critique. Science and Public Policy, 8(2):136– 141, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1981:QA

[Gol81b] Maurice Goldsmith. Quest, an autobiography. Science and Public Policy, 8(3):252–253, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1982:SSC

[Gol82] Dr Maurice Goldsmith. Science for social change. Science and Public Policy, 9(3):114– 119, June 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1983:DJS

[Gol83] Maurice Goldsmith. Derek J.

de Solla Price (1922–1983). Science and Public Policy, 10(6):318, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1984:CIM

[Gol84a] Maurice Goldsmith. Communication, infection and the metacomputer. Science and Public Policy, 11(6):334, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1984:HF

[Gol84b] Maurice Goldsmith. Historians of the future. Science and Public Policy, 11(5):258, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1984:ST

[Gol84c] Maurice Goldsmith. ‘slavery’ through S and T. Science and Public Policy, 11(4):254–255, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1984:TSO

[Gol84d] Maurice Goldsmith. Tracing scientific origins. Science and Public Policy, 11(4):254, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1984:TAS

[Gol84e] Maurice Goldsmith. Twentieth anniversary sophistication. Science and Public Policy, 11(1):2, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldberg:1985:CC

[Gol85a] David J. Goldberg. Chomsky’s complaint. Science and Public Policy, 12(1):49–50, February 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1985:IIS

[Gol85b] Maurice Goldsmith. Individual impacts of science. Science and Public Policy, 12(4):225– 226, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1986:CST

[Gol86a] Maurice Goldsmith. Commonwealth science teachers. Science and Public Policy, 13(3): 124, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1986:CVE

[Gol86b] Maurice Goldsmith. Cultural values and economic development. Science and Public Policy, 13(1): 2, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1986:NDW

[Gol86c] Maurice Goldsmith. New definitions of wealth and productivity. Science and Public Policy, 13(4): 184, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1987:TJN

[Gol87a] Dr Maurice Goldsmith. Tribute to Joseph Needham. Science and Public Policy, 14(2):107– 108, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1987:SS

[Gol87b] Maurice Goldsmith. The science of science. Science and Public Policy, 14(1):12–18, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1987:TY

[Gol87c] Maurice Goldsmith. Twenty years on. Science and Public Policy, 14(1):2, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1988:TNE

[Gol88] Maurice Goldsmith. Toward a new era for science. Science and Public Policy, 15(3):130, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goodin:1980:ESR

[Goo80] Robert E. Goodin. The ethics of social risks. Science and Public Policy, 7(5):318–327, October 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goreish:1984:PCI

[Gor84] M. Goreish. Policies for coping with the introduction of new technologies into the Arab region. Science and Public Policy, 11(2):83– 85, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gowing:1981:SSS

[Gow81] Margaret Gowing. Science, space and survival. Science and Public Policy, 8(2):155–156, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Glaeser:1983:CIC

[GPH83] Bernhard Glaeser and Kevin Phillips-Howard. Comparative investigation of conceptual models in human ecology. Science and Public Policy, 10(1):10–20, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gingras:1988:ERP

[GR88] Yves Gingras and Jacques Rivard. Energy R&D policy in Canada. Science and Public Policy, 15(1): 35–42, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Green:1982:ICB

[Gre82a] David Green. Innovation, conservation and Britain: some implications of the world problematic. Science and Public Policy, 9(3): 156–162, June 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Green:1982:SS

[Gre82b] David Green. Stratospheric strife. Science and Public Policy, 9(5): 274–275, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Green:1983:SBP

[Gre83] David Green. Shaping Britain’s population growth. Science and Public Policy, 10(2):62–64, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gregory:1985:PIG

[Gre85] Geoff Gregory. Performance indicators for a government science department. Science and Public Policy, 12(3):107–112, June 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Green:1986:NTG

[Gre86] Ken Green. New technology and global living standards. Science and Public Policy, 13(5):307–308, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Green:1987:BRU

[Gre87] Dr Kenneth Green. British R&D under pressure. Science and Public Policy, 14(4):241–242, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Griffiths:1980:AER

[Gri80] Richard F. Griffiths. Acceptability and estimation in risk management. Science and Public Policy, 7(3):154–161, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Griffith:1984:LHC

[Gri84] W. P. Griffith. Light on the heat: The CO_{2} question. Science and Public Policy, 11(1):60– 61, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Groenewegen:1989:CRI

[Gro89] Peter Groenewegen. Chemical risk issues. Science and Public Policy, 16(2):132–133, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1985:TH

[GS85] Maurice Goldsmith and Stanley Segal. Technology and the handicapped. Science and Public Policy, 12(6):324–326, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1980:SPS

[Gum80] Dr Philip Gummett. Scientists in the public service. Science and Public Policy, 7(6):419–432, December 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1981:HME

[Gum81a] Philip Gummett. How much is enough? The MX, Cruise and Trident decisions. Science and Public Policy, 8(5):376–388, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1981:SSP

[Gum81b] Philip Gummett. The state of science policy in the UK. Science and Public Policy, 8(2):99– 109, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1981:TLL

[Gum81c] Philip Gummett. Too late for light? Science and Public Policy, 8(5):410–411, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1982:CR

[Gum82a] Philip Gummett. Contentious reading. Science and Public Policy, 9(3):168, June 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1982:SSI

[Gum82b] Philip Gummett. Supply-side insights. Science and Public Policy, 9(5):267, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1984:LW

[Gum84a] Dr Philip Gummett. A look at the world. Science and Public Policy, 11(4):252, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1984:STA

[Gum84b] Dr Philip Gummett. Second time around. Science and Public Policy, 11(4):251–252, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1985:MB

[Gum85] Philip Gummett. Meanwhile, back in 1945. Science and Public Policy, 12(6):353–354, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1986:HIG

[Gum86a] Philip Gummett. How to influence government policy. Science and Public Policy, 13(1):60– 61, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1986:OS

[Gum86b] Philip Gummett. Official secrets. Science and Public Policy, 13(2): 115–116, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1987:EBB

[Gum87a] Dr Philip Gummett. Europe’s bep, bap etc. Science and Public Policy, 14(5):304–305, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1987:DMT

[Gum87b] Philip Gummett. Diversity in military technology. Science and Public Policy, 14(6):357–360, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gummett:1989:WP

[Gum89] Dr Philip Gummett. Weapons in peacetime. Science and Public Policy, 16(1):57–59, February 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gunawardena:1980:DCF

[Gun80] Wimal Gunawardena. Developing countries and the frontiers of science and technology. Science and Public Policy, 7(3):207– 214, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gutmann:1980:ACR

[Gut80] Peter Gutmann. Assessing country risk. Science and Public Policy, 7(5):327–336, October 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goeller:1983:WDM

[GW83] H. E. Goeller and Alvin M. Weinberg. Where disarmers miss the point. Science and Public Policy, 10(6):317, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hagendijk:1982:SDN

[Hag82] R. P. Hagendijk. ‘science dynamics’ in The Netherlands. Science and Public Policy, 9(6):292–296, December 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hanna:1984:TCI

[Han84] Judith Hanna. Twentieth century inquisition. Science and Public Policy, 11(5):328, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Haque:1989:MPI

[Haq89a] M. Mainul Haque. A method for promoting indigenous innovations. Science and Public Policy, 16(1): 37–45, February 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Haque:1989:PII

[Haq89b] M. Mainul Haque. Promotion of indigenous innovations in the non-farm sector for achieving sustained rural development. Science and Public Policy, 16(5):299– 306, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Harrison:1982:R

[Har82] John Harrison. Remythologizing. Science and Public Policy, 9(5): 275–276, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Harrison:1983:BA

[Har83] John Harrison. Brandt again. Science and Public Policy, 10(5):263– 264, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Harwood:1986:NS

[Har86] Jonathan Harwood. Nazism selective? Science and Public Policy, 13(4):245–246, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hart:1987:SCM

[Har87] David Hart. Separating civil and military nuclear programmes in the UK and france; a feasibility study. Science and Public Policy, 14(4):189–198, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Harlin:1988:BE

[Här88a] Benedikt Härlin. Biotechnology in Europe. Science and Public Policy, 15(6):407–410, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Harwood:1988:MSA

[Har88b] Dr Jonathan Harwood. Making science accessible. Science and Public Policy, 15(2):125– 126, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Harwood:1989:EMP

[Har89] Dr Jonathan Harwood. Eugenics in the medical profession. Science and Public Policy, 16(2):133– 134, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hasted:1984:APR

[Has84] J. B. Hasted. Attitudes to psychical research. Science and Public Policy, 11(4):249–250, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hasegawa:1986:JWI

[Has86] Michiko Hasegawa. Japanese and Western ideas of internationalism. Science and Public Policy, 13(1): 16–19, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1983:GS

[Hay83] Georgy Hay. Giant steps. Science and Public Policy, 10(5): 262, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1984:I

[Hay84a] George Hay. Introduction. Science and Public Policy, 11(5):259– 260, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1984:SCL

[Hay84b] George Hay. Sf concept listings. Science and Public Policy, 11(5): 284–326, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1985:NL

[Hay85] George Hay. A new language? Science and Public Policy, 12(2): 94–95, April 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1986:COa

[Hay86a] George Hay. Conscious from outside. Science and Public Policy, 13 (3):179, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1986:COb

[Hay86b] George Hay. Conscious from outside. Science and Public Policy, 13(5):303–304, October 1986.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1986:FSF

[Hay86c] George Hay. Feminist science fiction. Science and Public Policy, 13(4):246–247, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1987:HTU

[Hay87] George Hay. History turned upside down. Science and Public Policy, 14(4):240–241, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1988:PNS

[Hay88a] George Hay. Problem but no solution. Science and Public Policy, 15(1):69, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1988:SFI

[Hay88b] George Hay. Scientology of fundamental importance. Science and Public Policy, 15(3):184–187, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1989:CS

[Hay89a] George Hay. Conspiracy of silence.

Science and Public Policy, 16(4): 251–252, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1989:ECW

[Hay89b] George Hay. Everyone can win.

Science and Public Policy, 16(6):

373–374, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Harvey:1988:TEC

[HB88] Edward B. Harvey and John H. Blakely. Technology and employment in Canada. Science and Public Policy, 15(1):43–56, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hellen:1981:E

[Hel81a] J. A. Hellen. Ecocatastrophe. Science and Public Policy, 8(4):336– 339, August 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hellen:1981:EN

[Hel81b] J. A. Hellen. Exploding numbers. Science and Public Policy, 8(5): 413–414, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hellen:1981:US

[Hel81c] J. A. Hellen. The United States in the 1980s. Science and Public Policy, 8(3):247–249, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hellen:1981:DCP

[Hel81d] J. A. Hellen, MA Oxon, Dr Phil Bonn. Demographic change and public policy in Egypt and Nepal: some long-term implications for development planning. Science and Public Policy, 8(4): 308–336, August 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hellen:1982:LL

[Hel82a] J. A. Hellen. Law and lore. Science and Public Policy, 9(6):333–334, December 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hellen:1982:PPN

[Hel82b] J. A. Hellen. Preventive prescription, no posset. Science and Public Policy, 9(5):269–273, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hellen:1983:EH

[Hel83a] J. A. Hellen. Engineering health.

Science and Public Policy, 10(5): 256, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hellen:1983:SIL

[Hel83b] J. A. Hellen. Sitz im leben J. A. Hellen: Common Landscape of America, 1580–1845. Science and Public Policy, 10(4):204– 205, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hesketh:1982:ECE

[Hes82a] R. V. Hesketh. European churches and the energy issue: Official statements, reports, comments, 1975–1979; nuclear power and public policy: The social and ethical problems of fission technology; the nuclear barons; nuclear energy and the environment environmental sciences and applications, volume II. Science and Public Policy, 9(4):219–220, August 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hesketh:1982:ECP

[Hes82b] R. V. Hesketh. The export of civil plutonium. Science and Public Policy, 9(2):64–70, April 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hill:1981:QS

[Hil81] Peter Hill. Questions and speculations. Science and Public Policy, 8(5):408–409, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hills:1984:POF

[Hil84a] P. J. Hills. Book review: Planning for the Office of the Future, Planning for Electronic Mail, Planning for Word Processing, Planning for Telecommunications, Planning for Office Microcomputers and Planning for Information Handling all edited by Alan J. Simpson, Gower, Aldershot, UK. Science and Public Policy, 11(6): 394, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hills:1984:AHC

[Hil84b] Philip Hills. Aspects of human communication. Science and Public Policy, 11(6):381, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hills:1984:HCA

[Hil84c] Philip Hills. Human communication in an age of electronic revolution. Science and Public Policy, 11 (6):335–337, December 1984. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hilgartner:1985:HSC

[Hil85] Stephen Hilgartner. How safe is the car? Science and Public Policy, 12(6):357–358, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hill:1986:ECT

[Hil86a] Stephen Hill. Eighteen cases of technology transfer to Asia/ Pacific region countries. Science and Public Policy, 13(3):162– 169, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hill:1986:DLS

[Hil86b] Stephen Hill. From dark to light: seeing development strategies through the eyes of S&T indicators. Science and Public Policy, 13(5):275–284, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hinsley:1981:RDA

[Hin81] F. H. Hinsley. Reflections on the debate about the nuclear weapons. Science and Public Policy, 8(3): 174–180, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holdsworth:1988:IEN

[HL88] Dick Holdsworth and Gordon Lake. Integrating Europe: the new R&D calculus. Science and Public Policy, 15(6):411–425, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hochstrasser:1986:DES

[Hoc86] U. Hochstrasser. Development of a European science policy. Science and Public Policy, 13(5):265– 268, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hodgson:1983:FNP

[Hod83] P. E. Hodgson. The future of nuclear power. Science and Public Policy, 10(1):2–9, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hoong:1986:SAS

[Hoo86] Ang Wai Hoong. Out-of-school activities in science and technology in Singapore. Science and Public Policy, 13(3):155–161, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Horn:1980:ECE

[Hor80] B. Ray Horn. Environmental communication and education needs. Science and Public Policy, 7(3): 175–185, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hourani:1982:TPE

[Hou82] Benjamin T. Hourani. Teilhard’s political ecumene: Empire or Commonwealth? Science and Public Policy, 9(6):313–323, December 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hourani:1983:TCW

[Hou83] Benjamin T. Hourani. Toward a Commonwealth: a win-win strategy for US business. Science and Public Policy, 10(4):198–201, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hourani:1987:TCO

[Hou87] Benjamin T. Hourani. Towards the 21st century: the organization of power in post-industrial society. Science and Public Policy, 14(4): 217–229, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hay:1985:AA

[HR85] George Hay and Dr John Radford. Anomalies and absurdities. Science and Public Policy, 12(3): 106, June 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hryniszak:1981:ECO

[Hry81] W. Hryniszak. Engineering cooperation. Science and Public Policy, 8(3):227–238, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hughes:1983:ECT

[Hug83] Trevor C. Hughes. Evaporation control in the Third World. Science and Public Policy, 10(6):295– 300, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Idso:1984:SPP

[Ids84] S. B. Idso. Science and public policy implications of the greenhouse effect. Science and Public Policy, 11(3):178–179, June 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ii:1982:USP

[Ii82] George A. Keyworth Ii. US science policy and the need for excellence. Science and Public Policy, 9(1):2– 5, February 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ingelstam:1987:TSC

[Ing87] Lars Ingelstam. Technology and social change in Linköping. Science and Public Policy, 14(5):297– 298, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irigoien:1985:PEM

[Iri85] Josu Irigoien. The practice and experience of the Mondragon Basque co-operatives. Science and Public Policy, 12(2):72–78, April 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1984:SS

[Irw84] Dr A. Irwin. Science and society.

Science and Public Policy, 11(4): 253, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1985:CE

[Irw85] Alan Irwin. The corporate earthquake. Science and Public Policy, 12(4):227–228, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1986:APB

[Irw86] Alan Irwin. Applied philosophy at its best. Science and Public Policy, 13(4):247–248, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1987:TR

[Irw87] Alan Irwin. Technological risk.

Science and Public Policy, 14(4): 242–244, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1988:GR

[Irw88a] Alan Irwin. ‘greening’ research.

Science and Public Policy, 15(2): 124–125, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1988:TM

[Irw88b] Alan Irwin. Thoughtful and meticulous. Science and Public Policy, 15(5):360–361, October 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1988:VLP

[Irw88c] Alan Irwin. Voices of the local people. Science and Public Policy, 15(3):181–182, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Irwin:1989:EPA

[Irw89] Alan Irwin. Environmental policy and the academic community. Science and Public Policy, 16(4):252– 253, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jacobsson:1980:TCE

[Jac80]  Staffan Jacobsson.  Technical

change: employment and distribution in LDCs. Science and Public Policy, 7(2):124–135, April 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jackson:1987:CHP

[Jac87] Dr Timothy Jackson. Combined heat and power — a test-case for small-scale private generation in Britain. Science and Public Policy, 14(5):263–276, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jahoda:1986:AIO

[Jah86] Marie Jahoda. Artificial intelligence: an outsider’s perspective. Science and Public Policy, 13(6): 333–340, December 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jasinski:1988:PIP

[Jas88] Andrzei H. Jasinski. Public innovation policy measures and the process of economic reform in Poland. Science and Public Policy, 15(3):137–143, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jasinski:1989:BEP

[Jas89] Dr Andrzej H. Jasinski. Is the British experience with public innovation policy relevant in a centrally planned economy? Science and Public Policy, 16(1):27– 36, February 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jenkins:1982:CTC

[Jen82a] Norman Jenkins. Controlling technological change. Science and Public Policy, 9(2):103, April 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jenkins:1982:UNC

[Jen82b] Norman Jenkins. The unjustified neglect of combined heat and power. Science and Public Policy, 9(1):28–34, February 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jarrett-Kerr:1981:ELS

[JK81] Rev Fr Martin Jarrett-Kerr. The extreme of ludicrosity: some theological options. Science and Public Policy, 8(5):396–401, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jarrett-Kerr:1982:DG

[JK82] Martin Jarrett-Kerr. Deference to genes. Science and Public Policy, 9(2):100–103, April 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Johnston:1986:CTS

[Joh86] Ron Johnston. The centre for technology and social change. Science and Public Policy, 13(2):113– 114, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jones:1981:BRB

[Jon81a] Peter M. S. Jones. Book review: Scientists in Whitehall, by Philip Gummett Manchester University Press, 1980. Science and Public Policy, 8(2):163–164, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jones:1981:LP

[Jon81b] R. Russell Jones. Lead pollution. Science and Public Policy, 8(3): 195–202, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jones:1985:PRW

[Jon85a] Derek C. Jones. The potential role of workers’ co-operatives in Poland. Science and Public Policy, 12(2):61–66, April 1985. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jones:1985:SAT

[Jon85b] J. V. S. Jones. Strategic aspects of technology and science policy in Africa. Science and Public Policy, 12(4):175–190, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jordan:1986:BIH

[Jor86] Bill Jordan. Basic incomes and hidden economic potential. Science and Public Policy, 13(4):218– 220, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Joseph:1988:PDH

[Jos88] R. A. Joseph. The politics of defining high technology in Australia. Science and Public Policy, 15(5): 343–353, October 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Joseph:1989:SVM

[Jos89] R. A. Joseph. Silicon Valley myth and the origins of technology parks in Australia. Science and Public Policy, 16(6):353–365, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jakeman:1983:ARA

[JS83] A. J. Jakeman and Rodney W.

Simpson. The application of a risk assessment strategy to determine air pollution modeling and monitoring policy. Science and Public Policy, 10(6):289–294, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kapur:1989:WWC

[Kap89] C. Kapur. World-wide cooperation. Science and Public Policy, 16(5):311–313, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Keck:1984:MA

[Kec84] Dr Otto Keck. Not by money alone …. Science and Public Policy, 11(1):58–60, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kellerman:1982:CG

[Kel82] Dr E. W. Kellerman. Can S + T = growth? Science and Public Policy, 9(5):267–269, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kellerman:1984:PP

[Kel84] Dr E. W. Kellerman. Poor prospects. Science and Public Policy, 11(1):65–66, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kellerman:1987:DNE

[Kel87a] E. W. Kellerman. Dawn of a new era. Science and Public Policy, 14 (3):179–180, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kelly:1987:WO

[Kel87b] Dr Alison Kelly. Wasted opportunity. Science and Public Policy, 14 (6):352–353, December 1987. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kemp:1987:WSPa

[Kem87a] Stewart Kemp. Is the world a safe place? Science and Public Policy, 14(2):111–112, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kemp:1987:WSPb

[Kem87b] Stewart Kemp. Is the world a safe place? Science and Public Policy, 14(5):305–306, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Khan:1987:IMI

[Kha87] M. U. Khan. Impact of microelectronics in India. Science and Public Policy, 14(4):199–206, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kibble:1981:CP

[Kib81] T. W. B. Kibble. Case not proved.

Science and Public Policy, 8(6): 490–492, December 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kinnock:1985:PSW

[Kin85] Neil Kinnock. Putting science to work: a strategy for the future. Science and Public Policy, 12(6): 299–307, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

King:1987:CSC

[Kin87]  Alexander King.  Common-

wealth scientific co-operation during World War II. Science and Public Policy, 14(3):151–154, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kirschenbaum:1985:CIE

[Kir85] Alan Kirschenbaum. Computerization: impact on employment for the disabled. Science and Public Policy, 12(6):331–336, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kamenetzky:1989:AHN

[KM89a] Mario Kamenetzky and Robert H. Maybury. Agriculture in harmony with nature. Science and Public Policy, 16(2):73–82, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kamenetzky:1989:CE

[KM89b] Mario Kamenetzky and Robert H. Maybury. Contribution of education. Science and Public Policy, 16 (2):107–118, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kamenetzky:1989:ESE

[KM89c] Mario Kamenetzky and Robert H. Maybury. Ensuring a sustainable energy future. Science and Public Policy, 16(2):91–106, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kamenetzky:1989:MTN

[KM89d] Mario Kamenetzky and Robert H. Maybury. Management of technology and natural resources. Science and Public Policy, 16(2):66– 67, April 1989. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Kamenetzky:1989:TFR

[KM89e] Mario Kamenetzky and Robert H.

Maybury. Technical and financial rationality with concern for human needs and ecosystems: a conceptual framework. Science and Public Policy, 16(2):68–72, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kamenetzky:1989:TIN

[KM89f] Mario Kamenetzky and Robert H.

Maybury. Technology, industry and natural resources. Science and Public Policy, 16(2):83– 90, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kodama:1986:JIM

[Kod86] Fumio Kodama. Japanese innovation in mechatronics technology. Science and Public Policy, 13(1): 44–51, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Koorna:1985:FSU

[Köö85] Arno Arturovitch Köörna. Is fundamental science under threat in Great Britain? Science and Public Policy, 12(2):95–97, April 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Krull:1984:BRB

[Kru84a] Robert Krull. Book review: The Elements of Friendly Software Design by Paul Heckel, Warner, New York, NY, USA, 1982. Science and Public Policy, 11(6):394, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Krull:1984:VDE

[Kru84b] Robert Krull. Visual design of educational computer programs: Lessons from children’s TV. Science and Public Policy, 11(6):386– 387, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Koptas:1987:PAT

[KW87] Grazyna Koptas and Wieslaw Wisniewski. Public attitudes towards scientific knowledge in Poland. Science and Public Policy, 14(3):118–132, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lake:1981:PSS

[Lak81] Gordon J. Lake. The potential of the STSA support centre. Science and Public Policy, 8(1):45– 46, February 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lalor:1984:PUE

[Lal84] G. C. Lalor. Possibilities for university education in an age of electronic revolution. Science and Public Policy, 11(6):374–377, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lamm:1984:IHO

[Lam84] Zvi Lamm. Ideologies in a hierarchical order: a neglected theory. Science and Public Policy, 11(1): 40–46, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langford:1984:SS

[Lan84] David Langford. Science out of sf?

Science and Public Policy, 11(5): 273–274, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Landry:1989:BEM

[Lan89] Réjean Landry. Barriers to efficient monitoring of science, technology and innovation through public policy. Science and Public Policy, 16(6):345–352, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lassere:1980:CLE

[Las80] Pierre Lassere. The contribution of lagoons to economic growth. Science and Public Policy, 7(4): 288–292, August 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lederman:1985:STE

[Led85] Leonard L. Lederman. Science and technology in Europe: a survey. Science and Public Policy, 12(3): 131–143, June 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Leiss:1988:ITP

[Lei88] William Leiss. Industry, technology and the political agenda in Canada; the case of government support for R&D. Science and Public Policy, 15(1):57– 65, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lensen:1988:DIS

[Len88] Anton Lensen. Designing the information society. Science and Public Policy, 15(3):182–183, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Leydesdorff:1988:PFB

[Ley88a] Loet Leydesdorff. Performance figures for British science. Science and Public Policy, 15(4): 270, August 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Leydesdorff:1988:PMN

[Ley88b] Loet Leydesdorff. Problems with the ‘measurement’ of national scientific performance. Science and Public Policy, 15(3):149–152, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lichfield:1982:CU

[Lic82]  Nathaniel Lichfield.  Contra-

dictions unresolved. Science and Public Policy, 9(5):278–279, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lichfield:1987:NAS

[Lic87] Nathaniel Lichfield. National approaches to science parks. Science and Public Policy, 14(5):301– 303, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lindh:1983:PTK

[Lin83] Gunnar Lindh. Problems in the transfer of knowledge about water resources and management. Science and Public Policy, 10(4):192– 197, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Liyanage:1987:LSL

[Liy87] Dr S. Liyanage. Lessons from Sri Lanka’s science and technology planning. Science and Public Policy, 14(2):59–70, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lasmon:1983:CPS

[LJ83] C. S. F. Lasmon and D. J. Jeffries. Current problems in sexually transmitted diseases. Science and Public Policy, 10(5):218– 223, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lederman:1986:RPS

[LLB86] Leonard L. Lederman, Rolf Lehming, and Jennifer S. Bond. Research policies and strategies in six countries: a comparative analysis. Science and Public Policy, 13(2):67–76, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lovelock:1981:PF

[Lov81] James E. Lovelock. A policy for fluorocarbons. Science and Public Policy, 8(3):203–205, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lowe:1983:WBA

[Low83] Ian Lowe. Who benefits from Australian energy research? Science and Public Policy, 10(6):278–284, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lucas:1984:UN

[Luc84] Arthur Lucas. Untaming nature. Science and Public Policy, 11(4): 250–251, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lumb:1982:UTR

[Lum82] Audrey E. Lumb. University teaching, research and libraries — countering the cuts in library spending. Science and Public Policy, 9(6):324–329, December 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lunan:1987:NTP

[Lun87a] Duncan Lunan. Notes towards a politics for survival. Science and Public Policy, 14(1):43–50, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lunan:1987:SFF

[Lun87b] Duncan Lunan. Science fact or fiction? Science and Public Policy, 14(1):54–56, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Macioti:1986:PES

[Mac86] Manfredo Macioti. Progress of European science and technology. Science and Public Policy, 13(2): 97–100, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Macioti:1989:IDS

[Mac89a] Manfredo Macioti. Industrial development and a sustainable environment. Science and Public Policy, 16(5):262–268, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Macioti:1989:VDV

[Mac89b] Manfredo Macioti. Varied and divergent views. Science and Public Policy, 16(5):314–315, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Magyar:1984:CPS

[Mag84] George Magyar. Changing pattern of sponsorship in physics. Science and Public Policy, 11(4):188– 190, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Maitland:1989:FIC

[Mai89] Donald Maitland. The future of international communications. Science and Public Policy, 16(5): 307–309, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Male:1987:HTS

[Mal87] Dr S. E. Male. High temperature superconductivity. Science and Public Policy, 14(6):362–364, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Manavu:1980:EA

[Man80a] R. M. Manavu. Energy in Africa. Science and Public Policy, 7(6): 476–478, December 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mansfield:1980:IUC

[Man80b] Edwin Mansfield. Innovation in the USA: 1. A change in the economic climate is required. Science and Public Policy, 7(3):161– 165, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Manasse:1982:WAH

[Man82] O. E. Manasse. Widening the access to homoeopathy. Science and Public Policy, 9(6):282–285, December 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marcum:1985:TLC

[Mar85] John M. Marcum. Technology leadership: co-operation, competition and interdependency. Science and Public Policy, 12(6):317–323, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Martin:1986:NS

[Mar86] Brian Martin. Nuclear suppression. Science and Public Policy, 13 (6):312–320, December 1986. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marchington:1987:LF

[Mar87a] Mick Marchington. Life in the future. Science and Public Policy, 14 (3):177–178, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marchington:1987:MUP

[Mar87b] Mick Marchington. Managementunion participation brewing. Science and Public Policy, 14(2):112– 113, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Martin:1988:NWS

[Mar88] Brian Martin. Nuclear winter:

science and politics. Science and Public Policy, 15(5):321–334, October 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Masuzoe:1986:JSO

[Mas86] Yoichi Masuzoe. Japanese society and the outside world. Science and Public Policy, 13(1):33– 37, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Matthews:1983:UPO

[Mat83] Paul Matthews. The UK post office, British telecom and the law. Science and Public Policy, 10(6):

304–306, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mayne:1987:FGS

[May87] Alan J. Mayne. Fifth generation science policies. Science and Public Policy, 14(5):299–300, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mayor:1988:NDU

[May88] Federico Mayor. A new design for Unesco. Science and Public Policy, 15(6):426–430, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mazumdar:1980:RFG

[Maz80] Subash Mazumdar. Realistic food goals for Africa. Science and Public Policy, 7(6):469–475, December 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McGovern:1989:FTT

[McG89] Fergus McGovern. The future of telecommunications technology. Science and Public Policy, 16(2): 129–130, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meadows:1986:ASC

[Mea86] A. J. Meadows. Amateur science and communication. Science and Public Policy, 13(5):285– 289, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Medawar:1984:DMM

[Med84] Sir Peter Medawar. Does mind matter? Science and Public Policy, 11(4):246–248, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mellanby:1986:DO

[Mel86] Jean Mellanby. Disabled by obsessions. Science and Public Policy, 13(4):249, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Menon:1980:RG

[Men80] M. G. K. Menon. Rules of the game. Science and Public Policy, 7 (1):9–20, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mepham:1989:SPB

[Mep89] Dr T. B. Mepham. Science and the politics of breastfeeding: birthright or birth rite? Science and Public Policy, 16(3):181– 191, June 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meyers:1982:BSH

[Mey82] Aaron L. Meyers. Born secret: The H-bomb. the progressive case and national security. Science and Public Policy, 9(4):217– 219, August 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mg:1981:LG

[Mg81] Mg. Limited guide. Science and Public Policy, 8(2):165, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mg:1986:A

[Mg86] Mg. Announcement. Science and Public Policy, 13(4):249, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mlawa:1989:FG

[Mla89] Hasa Mfaume Mlawa. Filling a gap. Science and Public Policy, 16 (3):199–200, June 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mohan:1986:WPE

[Moh86] Dinesh Mohan. West patronising East. Science and Public Policy, 13(1):61–62, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Moore:1986:ACP

[Moo86] J. D. L. Moore. Afro-centric perspective. Science and Public Policy, 13(2):118–119, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Moore:1989:AP

[Moo89] John Moore. Arts of peace. Science and Public Policy, 16(3):192– 196, June 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Morehouse:1980:SBC

[Mor80] Ward Morehouse. Surviving the Brandt commission. Science and Public Policy, 7(6):483–485, December 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Moravcsik:1981:STA

[Mor81a] Michael J. Moravcsik. Scientific and technological assistance and the challenge of the 80’s. Science and Public Policy, 8(1):40– 44, February 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Morehouse:1981:TEB

[Mor81b] Ward Morehouse. Technology and equity in black holes: the ‘refraction effect’ of technology on social change. Science and Public Policy, 8(6):487–489, December 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Morphet:1981:PPA

[Mor81c] C. Morphet. Positivist and political approaches to interdisciplinarity. Science and Public Policy, 8 (1):18–22, February 1981. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Morgan:1984:FCP

[Mor84] Chris Morgan. Following the classic pattern. Science and Public Policy, 11(5):330, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Morehouse:1985:RBS

[Mor85] Ward Morehouse. Regional benefits from serious thought? Science and Public Policy, 12(6):354–355, December 1985. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Morris:1986:CBP

[Mor86] Colin Morris. Challenge of British purity. Science and Public Policy, 13(5):301, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McBride:1981:WER

[MS81] J. R. McBride and I. R. Sanderson. The work of the energy research group of the open University. Science and Public Policy, 8(1):13–17, February 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mullin:1981:SPN

[Mul81] James Mullin. Science policy in its new context. Science and Public Policy, 8(2):??, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mullins:1987:ERP

[Mul87] Nicholas C. Mullins. Evaluating research programs: measurement and data sources. Science and Public Policy, 14(2):91– 98, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Munk:1981:GIT

[Mun81] Anders Munk. Getting it together. Science and Public Policy, 8(5): 418, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Munk:1982:RI

[Mun82] Anders Munk. Rationalized intuition. Science and Public Policy, 9(5):277–278, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Munroe:1983:DD

[Mun83] Pat Munroe. Dirge on dirigisme.

Science and Public Policy, 10(6): 315–316, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Murphy:1985:MVT

[Mur85] John W. Murphy. A modern view on the transfer of technology. Science and Public Policy, 12(3):144– 148, June 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Murawski:1986:LYN

[Mur86a] Krzysztof Murawski. Love your neighbour. Science and Public Policy, 13(4):244, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Murphy:1986:GP

[Mur86b] Caroline Murphy. A glimpse of the past. Science and Public Policy, 13(5):305–306, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mushakoji:1980:TCD

[Mus80] Kinhide Mushakoji. Tradition and change in developing countries. Science and Public Policy, 7(1): 20–30, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nagaich:1982:SDP

[Nag82] B. B. Nagaich. The storage and dehydration of potato to feed growing populations. Science and Public Policy, 9(4):210– 216, August 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nance:1982:IWW

[Nan82] C. T. Nance. Inaction on windships: wisdom or folly? Science and Public Policy, 9(2):88– 96, April 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Napolitano:1988:ETC

[Nap88] Giovanni Napolitano. European technological co-operation: the Italian participation in Eureka. Science and Public Policy, 15(6): 376–382, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Napolitano:1989:CP

[Nap89] Giovanni Napolitano. Collaborative projects. Science and Public Policy, 16(6):378–379, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Narin:1987:BTE

[Nar87] Francis Narin. Bibliometric techniques in the evaluation of research programs. Science and Public Policy, 14(2):99–106, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Narjes:1988:ETC

[Nar88] Karl-Heinz Narjes. Europe’s technological challenge: a view from the European commission. Science and Public Policy, 15(6):395–402, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

NDoye:1980:NQP

[N’D80] Thianar N’Doye. Nutrition is a question of philosophy. Science and Public Policy, 7(4):292– 299, August 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ndebbio:1985:GPE

[Nde85] John E. Udo Ndebbio. Growth process and economic transformation through technological transfer in Africa. Science and Public Policy, 12(4):211–222, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nelson:1982:PPC

[Nel82] J. G. Nelson. Public participation in comprehensive resource and environmental management. Science and Public Policy, 9(5):240– 250, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Neuburger:1985:FJF

[Neu85a] Henry Neuburger. Four of Japan’s five largest. Science and Public Policy, 12(4):229–230, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Neuburger:1985:WTD

[Neu85b] Henry Neuburger. Wave to technological determinism. Science and Public Policy, 12(5): 290, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Neuberger:1986:SB

[Neu86] Henry Neuberger. Standing back. Science and Public Policy, 13(2): 116–117, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nelson:1982:LSI

[NF82] J. G. Nelson and Terry Fenge. Large scale industrial development in rural areas — a humanecological perspective. Science and Public Policy, 9(3):144– 155, June 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nixon:1987:SC

[Nix87] Sir Edwin Nixon. Surviving in the 21st century. Science and Public Policy, 14(1):3–11, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Narin:1988:LAM

[NSA^{+}88] Francis Narin, Kimberly Stevens, Joe Anderson, Peter Collins, John Irvine, Phoebe Isard, and Ben Martin. On-line approaches to measuring national scientific output: a cautionary tale. Science and Public Policy, 15(3):153– 161, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Oakley:1984:CCA

[Oak84] Brian Oakley. Computers and cooperation: The Alvey programme of research in information technology. Science and Public Policy, 11 (6):360–363, December 1984. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Okeke:1986:AWS

[Oke86] Eunice A. C. Okeke. Attracting women into science-based occupations: problems and prospects. Science and Public Policy, 13(3): 147–154, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Oldham:1989:TOD

[Old89] Geoffrey Oldham. Tibet opens its doors to foreign scientists. Science and Public Policy, 16(2):126– 128, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ohene-Manu:1987:TDS

[OM87] J. Ohene-Manu. Technology and development: socio-economic implications for Ghana. Science and Public Policy, 14(1):31–36, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

OToole:1983:BRB

[O’T83] J. J. M. O’Toole. Book review: Advice to a Young Scientist, P. B. Medawar, Harper & Row, Landon, 1980, 109 pp (cloth) £4.95. Science and Public Policy, 10(4): 215, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Palmer:1987:SIB

[Pal87] Arthur Palmer. Science, industry and a British parliamentary committee. Science and Public Policy, 14(3):145–150, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Munk:1981:MMN

[pAM81] Dr. phil Anders Munk. Methodical materialism: a necessary change in the concept of science. Science and Public Policy, 8(6): 479–486, December 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Munk:1983:R

[pAM83] Dr phil Anders Munk. Redundancy. Science and Public Policy, 10(2):84, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pavitt:1984:STI

[Pav84] Keith Pavitt. Science and technology indicators: eight conclusions. Science and Public Policy, 11(1): 21–24, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pavitt:1987:OTP

[Pav87] Keith Pavitt. The objectives of technology policy. Science and Public Policy, 14(4):182–188, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pearce:1981:NPP

[Pea81] David Pearce. Nuclear power and the public interest. Science and Public Policy, 8(1):23– 31, February 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Peng:1980:TW

[Pen80] Khor Kok Peng. The Third World.

Science and Public Policy, 7(6): 433–442, December 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Perera:1985:FAT

[Per85] Neomal N. Perera. A fresh approach to Third World self-help housing. Science and Public Policy, 12(5):286–287, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Petrella:1984:TEE

[Pet84] Riccardo Petrella. Technology and employment in Europe: Problems and proofs. Science and Public Policy, 11(6):352–359, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Phillips:1988:SSU

[Phi88] David Phillips. A strategy for science in the UK. Science and Public Policy, 15(1):3–12, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pick:1980:RDE

[Pic80] H. J. Pick. R and D for energy and raw-materials conservation. Science and Public Policy, 7(2):83– 98, April 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Piganiol:1986:IST

[Pig86] Pierre Piganiol. The interface of science and technology with society. Science and Public Policy, 13(5):259–264, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Peiting:1985:CUM

[PJ85] Yang Peiting and Liu Ji. The changing understanding and management of science in China. Science and Public Policy, 12(5):241– 252, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hellen:1983:TT

[pJAH83a] Dr phil J. A. Hellen. Trends or trendiness? Science and Public Policy, 10(2):79–81, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ma:1983:ADW

[pJAH83b] Dr phil J. A. Hellen, MA. Africa:

a drift or a wash? Science and Public Policy, 10(2):89– 91, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pollard:1982:PCS

[Pol82] John A. Pollard. Pre-conceptional sex selection. Science and Public Policy, 9(1):16–19, February 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Polya:1989:CC

[Pol89] Dr D. A. Polya. Colourful and comprehensive. Science and Public Policy, 16(4):255, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Poniatowski:1988:ETC

[Pon88] M. Poniatowski. Europe’s technological challenge: a view from the European parliament. Science and Public Policy, 15(6):383–394, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Porte:1984:BRB

[Por84a] Michael Porte. Book review:

Political Campaign Communications: Principles and Practices by Judith S. Trent and Robert V. Friedenberg, Praeger, New York, NY, USA, 1983. Science and Public Policy, 11(6):393–394, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Porte:1984:BRD

[Por84b] Michael Porte. Book review: Disconnecting Bell: The Impact of the AT&T Divestiture, edited by Harry M. Sooshan III, Pergamon, New York, NY, USA, 1984. Science and Public Policy, 11(6):393, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Postgate:1987:NFR

[Pos87] John Postgate. Nitrogen fixation research. Science and Public Policy, 14(4):236–237, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Power:1986:PHQ

[Pow86] Colin Power. Providing high quality teachers and support staff for the inter-related teaching of science, technology and mathematics. Science and Public Policy, 13 (3):134–146, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pritchard:1983:ESV

[Pri83a] Mark Pritchard. Energy: slanted views. Science and Public Policy, 10(4):210, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pritchard:1983:ST

[Pri83b] Mark Pritchard. Sun trap. Science and Public Policy, 10(5): 259, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Price:1985:EST

[Pri85] Geoffrey Price. Ethics in science and technology. Science and Public Policy, 12(4):228–229, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Price:1986:HP

[Pri86] Geoffrey Price. Historical perspective. Science and Public Policy, 13 (2):120–121, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Porte:1984:HCM

[PWG84] Michael Porte, Ernest Weiler, and Walter Gunn. Human communication by mass media of an imminent disaster. Science and Public Policy, 11(6):391–392, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pyke:1980:ED

[Pyk80] Magnus Pyke. Energy divine. Science and Public Policy, 7(1):65– 67, February 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pyke:1981:CES

[Pyk81] Dr Magnus Pyke. Cogito ergo sum. Science and Public Policy, 8 (1):5–12, February 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pyke:1987:DDA

[Pyk87] Dr Magnus Pyke. Dietary doctrines ancient and modern. Science and Public Policy, 14(6):347– 351, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1980:HLN

[Rah80] A. Rahman. History lessons neglected. Science and Public Policy, 7(4):303–304, August 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1985:EI

[Rah85a] A. Rahman. Economics in India. Science and Public Policy, 12(1): 51–52, February 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1985:JSIa

[Rah85b] A. Rahman. Jai Singh and the Indian renaissance. Science and Public Policy, 12(1):29–35, February 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1985:JSIb

[Rah85c] A. Rahman. Jai Singh and the Indian renaissance. Science and Public Policy, 12(2):88– 93, April 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1986:SCT

[Rah86] A. Rahman. Science and the cultural transformation: a history from India. Science and Public Policy, 13(5):290–294, October 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1987:IER

[Rah87] A. Rahman. Indian experience of renewable sources of energy. Science and Public Policy, 14(5):277– 285, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1988:STS

[Rah88] A. Rahman. Science, technology and spiritual values — a historical and social perspective. Science and Public Policy, 15(2):115– 118, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1989:SCD

[Rah89] A. Rahman. Science, culture and development. Science and Public Policy, 16(4):256–257, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ramesam:1987:STI

[Ram87] Dr V. Ramesam. Science and technology information programmes in India — some perspectives. Science and Public Policy, 14(1):37– 41, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ramanathan:1988:MTF

[Ram88a] Dr K. Ramanathan. Measurement of technology at the firm level. Science and Public Policy, 15(4):230– 248, August 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ramanathan:1988:ENS

[Ram88b] K. Ramanathan. Evaluating the national science and technology base: a case study on Sri Lanka. Science and Public Policy, 15(5): 304–320, October 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rao:1986:PMT

[Rao86] V. V. Subba Rao. Promotional measures for technology in Indian industry. Science and Public Policy, 13(2):106–112, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ravetz:1982:SFS

[Rav82] J. R. Ravetz. The social functions of science: a commemoration of J. D. Bernal’s vision. Science and Public Policy, 9(5):262– 266, October 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ray:1986:RAI

[Ray86] Tim Ray. Rare approach to industrial economics. Science and Public Policy, 13(3):181, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rosillo-Calle:1987:APF

[RCH87] F. Rosillo-Calle and J. Heaford. Alternatives to petroleum fuels for transport: Brazilian experience. Science and Public Policy, 14(6): 337–345, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Research:1980:AIE

[RD80] Atkins Research and Uk Development. Assessment of industrial environmental impact. Science and Public Policy, 7(2):145– 152, April 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rothwell:1989:TBS

[RD89] Roy Rothwell and Dr Mark Dodgson. Technology-based small and medium sized firms in Europe: The IRDAC results and their public policy implications. Science and Public Policy, 16(1):9– 18, February 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Raz:1987:TIA

[RDU87] Baruch Raz, Shabtay Dover, and Emily Udler. Technological innovations for agriculture in arid regions: Israel’s experience. Science and Public Policy, 14(4):207– 216, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McCausland:1984:GPH

[(Re84] Marcus McCausland (Ret’d). The gentle path to healing cancer: The Association for New Approaches to Cancer and its work. Science and Public Policy, 11(2):100– 105, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Read:1989:NMC

[Rea89] Nicholas Read. The ‘near market’ concept applied to UK agricultural research. Science and Public Policy, 16(4):233–238, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Russell:1980:AHC

[RF80] S. Russell and R. A. D. Ferguson. Assessing the health costs of fuel systems. Science and Public Policy, 7(5):365–376, October 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rabkin:1982:UUI

[RLH82] Yakov M. Rabkin and Jean-Jacques Lafitte-Houssat. USA– USSR interacademy exchanges. Science and Public Policy, 9(1):6– 11, February 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rothman:1980:HRD

[RM80] H. Rothman and R. D. Medford. Human resource and development problems. Science and Public Policy, 7(6):407–418, December 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Roberts:1981:STA

[Rob81a] Adam Roberts. Security through ‘alternative defence’. Science and Public Policy, 8(5):345–349, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Roberts:1981:PCL

[Rob81b] Colonel D. M. Roberts. Primary cancer of the liver. Science and Public Policy, 8(3):185– 190, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Robson:1983:POT

[Rob83] D. A. Robson. Peace in our time.

Science and Public Policy, 10(1): 44–45, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rogers:1981:AUP

[Rog81] C. S. Rogers. Academic unemployment: a personal view. Science and Public Policy, 8(6):458–464, December 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rosenbrock:1983:DAS

[Ros83] H. H. Rosenbrock. Designing automated systems — need skill be lost? Science and Public Policy, 10(6):274–277, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rothschild:1981:AIO

[Rot81] Emma Rothschild. Annex II, OECD report. Science and Public Policy, 8(2):141–143, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rothwell:1985:RTT

[Rot85] Dr Roy Rothwell. Reindustrialization and technology: Towards a national policy framework. Science and Public Policy, 12(3):113– 130, June 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rothwell:1986:SFC

[Rot86] Roy Rothwell. Small firms’ contributions to industrial innovation: small or large? Science and Public Policy, 13(3):170–172, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kronz:1984:PSC

[rpHK84] Dr rer pol Hermann Kronz, Dipl Ing. Patent statistics at the commission of the European communities. Science and Public Policy, 11(1):18–20, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1984:STA

[RT84] A. Rahman and S. Tirmizi. Science and technology archives: a guide to records. Science and Public Policy, 11(4):243–245, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ruberti:1988:SIE

[Rub88] Antonio Ruberti. Socialist issues in European science and technology. Science and Public Policy, 15 (6):403–406, December 1988. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rudig:1985:FEG

[Rüd85a] Wolfgang Rüdig. Future electricity generation. Science and Public Policy, 12(3):153–155, June 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rudig:1985:WRR

[Rüd85b] Wolfgang Rüdig. Weapons: the rulers and the ruled. Science and Public Policy, 12(5):291–293, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sabet:1980:APU

[Sab80] Adel A. Sabet. The Arab post-UNCSTD symposium. Science and Public Policy, 7(5):393–397, October 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Saha:1980:EER

[Sah80] Suranjit K. Saha. Environmental education for river-basin planning. Science and Public Policy, 7(4): 241–251, August 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sakamoto:1986:JEI

[Sak86] Masahiro Sakamoto. Japanese economy: an international perspective. Science and Public Policy, 13(1):38–43, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Salamon:1981:SPN

[Sal81] Jean-Jacques Salamon. Science policy in the new socio-economic context. Science and Public Policy, 8(2):86–91, April 1981. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sandford:1981:CCD

[San81]  Herbert A. Sandford.  Cur-

ing Colemanic depression. Science and Public Policy, 8(5):412, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sandford:1982:MMU

[San82] H. A. Sandford. Misleading maps of urbanization. Science and Public Policy, 9(1):12–15, February 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sapsed:1989:TUA

[Sap89] Jon Sapsed. Towards a usable anthropology. Science and Public Policy, 16(6):375–376, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sarett:1980:IUN

[Sar80] Lewis H. Sarett. Innovation in the USA: 3. new drugs are an endangered species. Science and Public Policy, 7(3):171–173, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Saviotti:1985:KD

[Sav85] Pier Paola Saviotti. Kaplinksi in dispute. Science and Public Policy, 12(6):355–357, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schiller:1984:NIT

[Sch84] Herbert Schiller. New information technologies and old objectives. Science and Public Policy, 11(6): 382–383, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Scott:1989:CTW

[Sco89] Pam Scott. Culling technological white elephants: Lessons from the Australian Animal Health Laboratory. Science and Public Policy, 16(1):47–51, February 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Segal:1987:PM

[Seg87] Stanley Segal. Practical and motivating. Science and Public Policy, 14(5):303–304, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Segal:1988:HTF

[Seg88] Howard P. Segal. High tech’s false nostalgia. Science and Public Policy, 15(3):189–190, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stolte-Heiskanen:1986:ESP

[SH86] V. Stolte-Heiskanen. Evaluation of scientific performance on the periphery. Science and Public Policy, 13(2):83–88, April 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sharma:1984:NEI

[Sha84] Dhirendra Sharma. Nuclear energy indicted. Science and Public Policy, 11(1):61–62, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sharma:1985:NPT

[Sha85] Dhirendra Sharma. Nuclear power and the Third World. Science and Public Policy, 12(3):155– 157, June 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sharif:1988:BTE

[Sha88a] Dr M. Nawaz Sharif. Basis for techno-economic policy analysis. Science and Public Policy, 15(4): 217–219, August 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sharif:1988:PIS

[Sha88b] Dr M. Nawaz Sharif. Problems, issues and strategies for S&T policy analysis. Science and Public Policy, 15(4):195–216, August 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shackley:1989:RRG

[Sha89] Simon J. Shackley. Regulation of the release of genetically manipulated organisms into the environment. Science and Public Policy, 16(4):211–223, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shirley:1984:SCE

[Shi84] Steve Shirley. Social consequences of the electronic revolution. Science and Public Policy, 11(6):350– 351, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shimada:1986:IIJ

[Shi86a] Haruo Shimada. International impact of Japan’s labor practices. Science and Public Policy, 13(1): 20–24, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shimo:1986:ITO

[Shi86b] Kotaro Shimo. Imported technology, originality and exports from Japan. Science and Public Policy, 13(1):3–6, February 1986. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shortland:1988:ELS

[Sho88] Dr Michael Shortland. Exemplary lives: a study of scientific autobiographies. Science and Public Policy, 15(3):170–179, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Simon:1980:IUC

[Sim80] Dorothy M. Simon. Innovation in the USA: 4. commercial applications should receive credit. Science and Public Policy, 7(3):173– 175, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Simpson:1987:NTS

[Sim87] Dr John Simpson. Nuclear test secrecy. Science and Public Policy, 14(2):114–115, April 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sinclair:1987:NSC

[Sin87a] Dr Craig Sinclair. Nato science Committee’s programme: promoting scientific mobility and international cohesion. Science and Public Policy, 14(3):169–170, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sinclair:1987:RSR

[Sin87b] Dr Craig Sinclair. Reflections on scientific research in the Soviet union. Science and Public Policy, 14(3):133–138, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sinclair:1988:CSP

[Sin88a] Craig Sinclair. Comparative science policies. Science and Public Policy, 15(6):433–434, December 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sinclair:1988:STG

[Sin88b] Craig Sinclair. Science and technology in Greece, Portugal and Turkey. Science and Public Policy, 15(5):354–356, October 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sinclair:1988:SNS

[Sin88c] Craig Sinclair. Science in the non-Soviet CMEA countries or Hamlet without the Prince? Science and Public Policy, 15(2):109– 113, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sirockin:1985:TWD

[Sir85] Geof Sirockin. The Third World a dumping ground? Science and Public Policy, 12(1):50–51, February 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Skolimowski:1986:WD

[Sko86]  Henryk Skolimowski.  White

death. Science and Public Policy, 13(4):243, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1985:CC

[Smi85a] George Telling Smith. Collectives for cures. Science and Public Policy, 12(4):231, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1985:CTD

[Smi85b] Michaela Y. Smith. Coming to terms with development. Science and Public Policy, 12(5):293– 294, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1986:PA

[Smi86] Denis Smith. Paternalistic authority. Science and Public Policy, 13 (4):245, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1987:ITT

[Smi87] Dr S. L. Smith. Information tecbnology: Taylorisation or human centred office systems. Science and Public Policy, 14(3):159– 167, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1988:QQ

[Smi88] George Teeling Smith. Quality not quantity. Science and Public Policy, 15(4):273, August 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:1989:AF

[Smi89] George Teeling Smith. Ahead of the field. Science and Public Policy, 16(5):315–316, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sorensen:1987:CSE

[Sør87] Bent Sørensen. Current status of energy supply technology and future requirements. Science and Public Policy, 14(5):252– 256, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sotiriou:1987:SPG

[Sot87] Dr D. S. Sotiriou. Science policy in Greece. Science and Public Policy, 14(1):23–30, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Spencer:1981:CW

[Spe81] Sir Kelvin Spencer. Changing the worldview. Science and Public Policy, 8(6):426–433, December 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Suzuki:1986:JNE

[SS86] Tatujiro Suzuki and Atsuyuki Suzuki. Japan’s nuclear energy policy. Science and Public Policy, 13(1):25–32, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schware:1985:RMT

[ST85] Robert Schware and Alice Trembour. Rethinking microcomputer technology transfer to Third World countries. Science and Public Policy, 12(1):15–20, February 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stansfield:1981:ORS

[Sta81] Ronald G. Stansfield. Operational research and sociology: a casestudy of cross-fertilizations in the growth of useful science. Science and Public Policy, 8(4):262– 280, August 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stansfield:1983:CC

[Sta83a] R. G. Stansfield. Cushioned crime.

Science and Public Policy, 10(2): 84–86, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stansfield:1983:SS

[Sta83b] Ronald G. Stansfield. Supernatural science. Science and Public Policy, 10(4):216, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stahel:1986:PLV

[Sta86a] Walter R. Stahel. Product life as a variable: the notion of utilization. Science and Public Policy, 13(4):185–193, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stahel:1986:RSS

[Sta86b] Walter R. Stahel. R&D in a sustainable society. Science and Public Policy, 13(4):196–203, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Staab:1988:TCR

[Sta88] Heinz A. Staab. Towards a climate of respect and trust in scientific research. Science and Public Policy, 15(3):144–148, June 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Staab:1989:DNF

[Sta89] Heinz A. Staab. The discovery of nuclear fission and the responsibility of the scientist. Science and Public Policy, 16(4):247– 250, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Steinhouse:1982:CTP

[Ste82] Herbert Steinhouse. Commercial television and the public alternative. Science and Public Policy, 9 (3):120–129, June 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stead:1984:CPS

[Ste84]  Humphrey Stead.  Canadian

patent statistics. Science and Public Policy, 11(1):10–17, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Steinberg:1985:IST

[Ste85] Gerald M. Steinberg. Israeli science and technology and the European Community goal. Science and Public Policy, 12(5):264– 268, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Strassman:1984:OEA

[Str84] Paul A. Strassman. Organizations in an electronic age. Science and Public Policy, 11(6):343–349, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stubbs:1988:RI

[Stu88] Peter Stubbs. From resignation to innovation. Science and Public Policy, 15(5):359–360, October 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sutcliffe:1980:SRB

[Sut80] J. R. Sutcliffe. Shipping risks at Braefoot Bay. Science and Public Policy, 7(5):356–364, October 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Swan:1986:LCG

[SW86] Simone Swan and Michaela Walsh. Local capital generation — rethinking employment initiatives. Science and Public Policy, 13(4): 207–208, August 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sweet:1986:HFC

[Swe86a] Colin Sweet. High fact content. Science and Public Policy, 13(1): 59–60, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sweet:1986:LDL

[Swe86b] Colin Sweet. Lies, damn lies and statistics. Science and Public Policy, 13(3):179–180, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sweet:1987:EA

[Swe87] Colin Sweet. Energy alternatives. Science and Public Policy, 14(5): 246–251, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Syrioti:1983:BRB

[Syr83] A. J. Syrioti. Book review: How to Repair, Reupholster and Refinish Furniture, Mel Marshall, Harper & Row, London, 1980 £6.95 cloth. Science and Public Policy, 10(4): 215–216, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tainsh:1981:STT

[Tai81a] A. Ramsay Tainsh. Sphinx thinks thring things. Science and Public Policy, 8(4):339–340, August 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tainsh:1981:MCD

[Tai81b] Maj A. Ramsay Tainsh. Migraine cured without drugs. Science and Public Policy, 8(3):250– 251, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tainsh:1983:AS

[Tai83a] A. Ramsay Tainsh. African socialism. Science and Public Policy, 10(2):81–82, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tainsh:1983:BND

[Tai83b] A. Ramsay Tainsh. The building of new delhi. Science and Public Policy, 10(1):45, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tainsh:1983:URA

[Tai83c] A. Ramsay Tainsh. The unhappy recipients of aid. Science and Public Policy, 10(1):38– 39, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tainsh:1984:RIR

[Tai84a] A. Ramsay Tainsh. Roots of the Iranian Revolution. Science and Public Policy, 11(1):65, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tainsh:1984:TP

[Tai84b] A. Ramsay Tainsh. Three pamphlets. Science and Public Policy, 11(1):64, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tal:1980:STP

[Tal80] Eliezer Tal. Science and technology policy in Israel in the 1970s. Science and Public Policy, 7(4): 275–281, August 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Talisayon:1983:NDG

[Tal83] Serafin D. Talisayon. New development goals and values in response to the global environmental crisis. Science and Public Policy, 10(1):21–26, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tal:1984:NSC

[Tal84] Eliezer Tal. The need for structural change in the science policy of Peru. Science and Public Policy, 11(2):77–82, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tal:1985:RC

[Tal85a] Eliezer Tal. R&D in Colombia. Science and Public Policy, 12(1): 3–14, February 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tal:1985:RP

[Tal85b] Eliezer Tal. R&D in Panama. Science and Public Policy, 12(5):253– 263, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tal:1989:STP

[Tal89] Dr Eliezer Tal. Science and technology policy advice in Israel. Science and Public Policy, 16(4):239– 246, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tame:1984:LLS

[Tam84] Chris R. Tame. Life, liberty and the stars. Science and Public Policy, 11(5):269–272, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tangermann:1981:FF

[Tan81] Stefan Tangermann. Food or famine. Science and Public Policy, 8(6):471–479, December 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tang:1986:CUT

[TC86] John C. S. Tang and Agnes C.

Cabrera. Computer usage in Thailand: a statistical survey. Science and Public Policy, 13(6):341–346, December 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thomas-Emeagwali:1989:TTE

[TE89] Gloria Thomas-Emeagwali. Technological transfer: early Europe and the third world. Science and Public Policy, 16(3):176–180, June 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thomas:1981:HES

[Tho81] M. Pugh Thomas. Human ecology and science and technology for development. Science and Public Policy, 8(1):47–54, February 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thomas:1982:OCV

[Tho82] M. Pugh Thomas. Onchocerciasis control in the Volta Basin. Science and Public Policy, 9(4):194– 202, August 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thompson:1983:SMR

[Tho83] Thomas W. Thompson. Schizophrenia — the myth and the reality. Science and Public Policy, 10(5): 251–252, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thomas:1985:LPN

[Tho85] Steve Thomas. Lost promise of nuclear power. Science and Public Policy, 12(6):348–350, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thring:1981:VE

[Thr81] M. W. Thring. Values in engineering. Science and Public Policy, 8(6):444–457, December 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thring:1983:ESJ

[Thr83a] M. W. Thring. Energy slaves and justice. Science and Public Policy, 10(6):316–317, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thring:1983:SEP

[Thr83b] Professor M. W. Thring. A scientific energy policy for Britain. Science and Public Policy, 10(3): 106–111, June 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thring:1984:GS

[Thr84] M. W. Thring. Getting up stream.

Science and Public Policy, 11(1): 66, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tigre:1983:PBC

[Tig83] Paulo Bastos Tigre. Protectionism in the Brazilian computer industry. Science and Public Policy, 10 (3):146–148, June 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Titanji:1985:BEV

[Tit85] Vincent Titanji. A bird’s eye view of problems of scientific research in Black Africa. Science and Public Policy, 12(4):207– 210, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tolley:1983:DHT

[Tol83] Rev Canon George Tolley. From disillusion to hope: a task for education. Science and Public Policy, 10(5):224–228, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tolley:1984:WR

[Tol84] George Tolley. Work and reality. Science and Public Policy, 11(2): 86–89, April 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Torrance:1982:TFK

[Tor82] Thomas F. Torrance. Transformation in the frame of knowledge. Science and Public Policy, 9(2): 97–100, April 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Toren:1984:NCS

[Tor84] Nina Toren. National cultures of science: a study of Soviet and American immigrant scientists in Israel. Science and Public Policy, 11(3):125–160, June 1984. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Trainer:1983:BRI

[Tra83a] F. E. Trainer. Brandt reports I and II — common mistake. Science and Public Policy, 10(6):285– 288, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Trainer:1983:WDM

[Tra83b] F. E. Trainer. Where disarmers miss the point. Science and Public Policy, 10(4):177–183, August 1983. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Trainer:1984:LE

[Tra84] Dr F. E. Trainer. Letter to the editor. Science and Public Policy, 11(1):54–56, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teeling-Smith:1980:STS

[TS80] G. Teeling-Smith. Socialist thesis on the spread of diseases. Science and Public Policy, 7(3):219– 220, June 1980. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teeling-Smith:1981:QS

[TS81a] G. Teeling-Smith. Qualified success. Science and Public Policy, 8(5):416–417, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teeling-Smith:1981:UW

[TS81b] G. Teeling-Smith. Unhealthy WHO? Science and Public Policy, 8(5):415–416, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teeling-Smith:1981:S

[TS81c] George Teeling-Smith. Synthesis. Science and Public Policy, 8 (2):148–151, April 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teeling-Smith:1983:CG

[TS83a] George Teeling-Smith. Counting germs. Science and Public Policy, 10(6):315, December 1983.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teeling-Smith:1983:DD

[TS83b] George Teeling-Smith. Drugs and drugs. Science and Public Policy, 10(2):82–83, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teeling-Smith:1983:NBA

[TS83c] George Teeling-Smith. NHS: belated assessment. Science and Public Policy, 10(5):256, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teeling-Smith:1984:SW

[TS84] G. Teeling-Smith. Sensationally wrong. Science and Public Policy, 11(1):62, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teeling-Smith:1986:SS

[TS86a] G. Teeling-Smith. Scottish sobriety. Science and Public Policy, 13 (3):180, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teeling-Smith:1986:HWR

[TS86b] George Teeling-Smith. Health as wealth: a response. Science and Public Policy, 13(6):362, December 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teeling-Smith:1989:LDL

[TS89]  George Teeling-Smith.  Lies,

damn lies and statistics. Science and Public Policy, 16(4):253– 254, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tsao:1989:CPF

[Tsa89] Jeffrey Y. Tsao. Consumer preferences and funding priorities in scientific research. Science and Public Policy, 16(5):294–298, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Uenohara:1986:JEI

[Uen86] Michiyuki Uenohara. Japanese electronics industry: from the 1950s into the information society. Science and Public Policy, 13 (1):7–15, February 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ulfvarson:1985:CHH

[Ulf85] Ulf Ulfvarson. Chemical health hazards in workplaces, prehistory to the present. Science and Public Policy, 12(1):21–28, February 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

VanRuskenveld:1987:TGB

[Van87] Yvonne Van Ruskenveld. Technology as good business. Science and Public Policy, 14(3):178– 179, June 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

VanRuskenveld:1988:PI

[Van88] Yvonne Van Ruskenveld. Partners in innovation? Science and Public Policy, 15(1):19–25, February 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vella:1981:END

[Vel81] Colonel Ethelwald E. Vella. Exotic new diseases. Science and Public Policy, 8(3):190–194, June 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vergragt:1987:LIL

[Ver87a] Dr Philip Vergragt. Like it or lump it? Science and Public Policy, 14(4):238–239, August 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vergragt:1987:OC

[Ver87b] Dr. Philip Vergragt. Only connect …. Science and Public Policy, 14 (6):354–355, December 1987. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Visser:1985:CCG

[Vis85] M. B. H. Visser. Commissions into controversies: genetic manipulation in The Netherlands. Science and Public Policy, 12(5):269– 272, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vanSteijn:1988:PTP

[vS88] Frans van Steijn. Part-time professors in The Netherlands. Science and Public Policy, 15(2):74– 80, April 1988. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wad:1985:STI

[Wad85] Atul Wad. Science, technology and industrialization in Africa. Science and Public Policy, 12(4): 164–174, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Walford:1983:BES

[Wal83a] George Walford. Beyond ecology: a study in systematic ideology. Science and Public Policy, 10(5):244–250, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Walford:1983:SIW

[Wal83b] George Walford. Systematic ideology: The work of Harold Walsby. Science and Public Policy, 10(1): 27–33, February 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ward:1985:RRT

[War85] David Ward. The rise and rise of transnational companies. Science and Public Policy, 12(2):84– 87, April 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Warboy:1986:US

[War86] Michael Warboy. Unfinished sympathy. Science and Public Policy, 13(3):176–177, June 1986. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Warren:1987:SMS

[War87] Andrew Warren. Saving megabucks by saving megawatts. Science and Public Policy, 14(5):287–291, October 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Warner:1989:ECN

[War89] Frederick Warner. Environmental consequences of nuclear war. Science and Public Policy, 16(1):53– 55, February 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watts:1982:CCE

[Wat82a] Michael Watts. Collapsing coexistence. Science and Public Policy, 9(5):277, October 1982. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watts:1982:MDS

[Wat82b] Michael Watts. Mass death of the soul. Science and Public Policy, 9 (6):333, December 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watts:1982:PDL

[Wat82c] Michael Watts. Prospects for development in Latin America. Science and Public Policy, 9(4):187– 193, August 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watts:1983:SA

[Wat83a] Lieut-Cdr Donald Watts. Strategic analysis. Science and Public Policy, 10(3):159–160, June 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watts:1983:RDW

[Wat83b] Lieut-Commander D. T. Watts.

The Royal Dockyards: waste of resources and the consequences. Science and Public Policy, 10(3):128– 133, June 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watts:1983:AR

[Wat83c] Michael Watts. Amazing race.

Science and Public Policy, 10(5): 258–259, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watts:1983:BPF

[Wat83d] Michael Watts. Brazil: planned futility. Science and Public Policy, 10(4):209–210, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watts:1983:LAO

[Wat83e] Michael Watts. Latin America:

origins, arguments and outcomes. Science and Public Policy, 10(4): 208–209, August 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watts:1983:LAP

[Wat83f] Michael Watts. Latin America: promises and promises. Science and Public Policy, 10(2):88– 89, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Watts:1983:SG

[Wat83g] Michael Watts. Source for a goose? Science and Public Policy, 10(5):257–258, October 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Webster:1989:PPS

[Web89] A. J. Webster. Privatisation of public sector research: The case of a plant breeding institute. Science and Public Policy, 16(4):224– 232, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Weholt:1984:HSS

[Weh84] Richard Weholt. The hard, the soft and the squishy. Science and Public Policy, 11(5):261–268, October 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Weinberg:1987:LS

[Wei87] Alvin M. Weinberg. Limits of science. Science and Public Policy, 14(1):51–52, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Weinberg:1989:UAT

[Wei89] Elliot H. Weinberg. US attitudes toward foreign science and technology. Science and Public Policy, 16(1):2–8, February 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wiener:1981:PAD

[Wie81] Aron Wiener. Paradigmatic assumptions for development and regulation in the 1980s. Science and Public Policy, 8(6):464–470, December 1981. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilson:1980:LCS

[Wil80] Andrew H. Wilson. Letter from Canada: Science policy lumbers on. Science and Public Policy, 7 (6):478–483, December 1980. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilson:1982:ICC

[Wil82] Andrew H. Wilson. Innovation in Canada: a commentary. Science and Public Policy, 9(6):302–312, December 1982. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilcock:1983:FG

[Wil83a] David Wilcock. Filling in the gaps.

Science and Public Policy, 10(2): 83–84, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilson:1983:LUP

[Wil83b] Rev Mervyn Wilson. Land use policy. Science and Public Policy, 10(3):158–159, June 1983. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Williams:1984:CRR

[Wil84a] Frederick Williams. The communications revolution revisited. Science and Public Policy, 11(6):338– 342, December 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilson:1984:CST

[Wil84b] Bryan J. Wilson. Communication in science, technology and mathematics education in the Commonwealth. Science and Public Policy, 11(4):191–200, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilson:1989:CUS

[Wil89a] Andrew H. Wilson. Canadian and US science and technology policy under free trade. Science and Public Policy, 16(2):119– 125, April 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilson:1989:GRA

[Wil89b] Andrew H. Wilson. The Gendron Report: Another view of Canadian science policy. Science and Public Policy, 16(5):269– 281, October 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wheale:1989:EH

[WM89] Peter Wheale and Ruth McNally. Environment and humanity. Science and Public Policy, 16(6):374– 375, December 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wolff:1985:TIS

[Wol85] Heinz Wolff. Technical innovations, social stigma and missed profits. Science and Public Policy, 12(6):327–330, December 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Worboys:1989:RSS

[Wor89] Dr Michael Worboys. Redundancy of science-society boundary. Science and Public Policy, 16(3):196– 198, June 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wright:1984:RHI

[Wri84] Robert R. Wright. Rationale, history and implications: The US experience. Science and Public Policy, 11(1):7–9, February 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilkinson:1983:MST

[WS83] Barry Wilkinson and Steven Smith. Management strategies for technical change. Science and Public Policy, 10(2):56–61, April 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilkinson:1985:EGE

[WW85] T. G. Wilkinson and P. Wilkinson.

Environmental groups and energy priorities. Science and Public Policy, 12(5):273–278, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Yapp:1983:PP

[Yap83] Brunsdon Yapp. Problems and proofs. Science and Public Policy, 10(6):307–309, December 1983. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Yoong:1986:MST

[Yoo86] Cheong Siew Yoong. Making science, technology and mathematics education relevant to youth in developing countries. Science and Public Policy, 13(3):125– 133, June 1986. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Yoxen:1985:BSM

[Yox85a] Dr Edward Yoxen. Broading students’ minds. Science and Public Policy, 12(4):230, August 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Yoxen:1985:PPM

[Yox85b] Edward Yoxen. Policy process and moral reasoning. Science and Public Policy, 12(5):288–289, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Yoxen:1987:BFA

[Yox87] Edward Yoxen. Biotechnology — four approaches. Science and Public Policy, 14(1):53–54, February 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ziman:1984:SU

[Zim84] John Ziman. Science down under. Science and Public Policy, 11(4): 253, August 1984. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ziman:1985:MSR

[Zim85] John Ziman. Modern social role of science. Science and Public Policy, 12(5):290–291, October 1985. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ziman:1989:FA

[Zim89] John Ziman. Fascinating and authoritative. Science and Public Policy, 16(4):258, August 1989. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zaal:1987:ASS

[ZL87] Rolf Zaal and Loet Leydesdorff.

Amsterdam science shop and its influence on university research: the effects of ten year of dealing with non-academic questions. Science and Public Policy, 14(6): 310–316, December 1987. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zuckerman:1981:OEA

[Zuc81] Arie J. Zuckerman. Observations on the enigma of aircraft cabin safety. Science and Public Policy, 8(5):402–404, October 1981. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).